### v1

# KNighter: Transforming Static Analysis with LLM-Synthesized Checkers

Abstract
Static analysis is a powerful technique for bug detection in critical systems like operating system kernels. However, designing and implementing static analyzers is challenging, time-consuming, and typically limited to predefined bug patterns. While large language models (LLMs) have shown promise for static analysis, directly applying them to scan large codebases remains impractical due to computational constraints and contextual limitations.
We present KNighter, the first approach that unlocks practical LLM-based static analysis by automatically synthesizing static analyzers from historical bug patterns. Rather than using LLMs to directly analyze massive codebases, our key insight is leveraging LLMs to generate specialized static analyzers guided by historical patch knowledge. KNighter implements this vision through a multi-stage synthesis pipeline that validates checker correctness against original patches and employs an automated refinement process to iteratively reduce false positives. Our evaluation on the Linux kernel demonstrates that KNighter generates high-precision checkers capable of detecting diverse bug patterns overlooked by existing human-written analyzers. To date, KNighter-synthesized checkers have discovered 70 new bugs/vulnerabilities in the Linux kernel, with 56 confirmed and 41 already fixed. 11 of these findings have been assigned CVE numbers. This work establishes an entirely new paradigm for scalable, reliable, and traceable LLM-based static analysis for real-world systems via checker synthesis.
1 Introduction
The reliability of fundamental systems, particularly operating system kernels, depends critically on effective defect detection techniques [23, 38, 12, 7, 21, 28, 35]. Static analysis [4] stands out as a powerful approach that can comprehensively examine source code without execution, enabling the detection of bugs in architecture-specific code, hardware-dependent drivers, and configurations that may be challenging to test in real environments [20, 21, 12, 28]. Unlike dynamic approaches such as fuzzing that require concrete execution environments and can only explore executed paths [2, 38, 9, 36], static analysis can theoretically examine all possible code paths, including those that are difficult or impossible to trigger during runtime. Plus, while formal verification [37, 14, 26] offers mathematical guarantees that static analysis cannot provide, it typically requires even more expertise and manual effort, making it impractical for large-scale systems like operating system kernels.
However, static analysis faces its own set of challenges in practice. Developing effective static analyzers requires deep expertise in both the target system and formal methods, along with substantial engineering effort to implement precise detection mechanisms [12, 7, 1]. This complexity often leads to static analyzers that focus on specific classes of bugs, potentially missing other critical defect patterns.
These observations highlight several key opportunities for advancing static analysis. First, the rich history of the fundamental systems, e.g., the Linux kernel, provides extensive examples of bug patterns that can inform static analysis approaches. Moreover, bug patterns identified in historical patch commits may still exist in the current codebase [12]. Third, beyond manually analyzing and summarizing bug patterns, modern large language models (LLMs) [3, 29, 11] have demonstrated remarkable capabilities in understanding both code and natural language descriptions.
Building on these insights, we propose leveraging LLMs to perform static analysis on real-world systems by learning from historical patch commits. However, directly applying LLMs to analyze real-world systems [38, 15, 18, 42, 10, 33] is challenging due to the enormous size of modern codebases—for example, the Linux kernel comprises over 30 million lines of code. The limited context windows of LLMs make it impractical to process an entire codebase at once, and the associated computational cost is prohibitive. In the case of the Linux kernel, which spans more than 300 million tokens, a single analysis pass using GPT-4o is estimated to cost approximately \$750. Moreover, LLMs are susceptible to hallucination issues, potentially leading to false positives [13, 10, 18]. Besides, LLMs also struggle to understand complex codebases [22].
Our solution addresses these challenges by using LLMs to synthesize static checkers based on historical bug patches rather than directly analyzing source code. These checkers, implemented using Clang Static Analyzer (CSA) [8], can then scan the kernel codebase effectively and efficiently. This approach offers several advantages:
  • It avoids the prohibitive costs and context limitations of processing large codebases directly with LLMs, thereby enhancing scalability.
  • The checkers can naturally evolve alongside kernel development without requiring constant LLM involvement.
  • We can validate the synthesized checkers against the original patches to mitigate hallucination concerns.
  • Our approach provides greater traceability and explainability compared to earlier techniques, as the synthesized checkers directly encode the bug patterns in a transparent, human-readable manner.
However, we observed that current LLMs struggle to synthesize and implement checkers end-to-end—a task that challenges even experienced researchers and developers with expertise in both static analysis and the target codebase. To address this limitation, we develop a multi-stage synthesis pipeline (§ 3.1) that breaks down the complex task into manageable steps. Furthermore, we design a fully automated refinement pipeline (§ 3.2) that iteratively improves the synthesized checkers to reduce false positives, ensuring their practical utility in real-world scenarios.
We implement our approach, KNighter, the first fully automated static analyzer generation pipeline. While our approach is highly generalizable to different software systems, we demonstrate its effectiveness by synthesizing checkers for 61 historical patch commits from the Linux kernel, successfully generating 37 plausible checkers. These checkers demonstrate high quality, with approximately 35% false positive rate. Our evaluation also presents that KNighter-detected vulnerabilities are orthogonal to existing expert-written static analyzers [1]. To date, checkers synthesized by KNighter have detected 70 new bugs/vulnerabilities in the Linux kernel, with 56 confirmed and 41 already fixed. Notably, 11 of these bugs have been assigned CVE numbers.
Our contributions are summarized below:
  • New paradigm. We introduce a pioneering LLM-based approach for synthesizing static analyzers from patch commits. To our knowledge, KNighter is the first fully automated static analyzer generation system, establishing a new paradigm for LLM-powered static analysis.
  • End-to-end pipeline. We implement KNighter as a comprehensive system featuring multi-stage synthesis and automated refinement pipelines for the Linux kernel. This design enables detection of diverse bug classes and adapts dynamically to emerging vulnerability patterns.
  • Systematic evaluation. We rigorously evaluate KNighter by synthesizing checkers for the patch commits spanning multiple bug categories. Our results demonstrate that KNighter generates high-precision checkers with practical false positive rates.
  • Real-world impact. KNighter-generated checkers have discovered 70 previously unknown bugs/vulnerabilities in the Linux kernel, with 56 confirmed by developers, 41 already fixed, and 11 assigned CVE numbers—validating both the approach’s effectiveness and practical security value.
Our code is available at ise-uiuc/KNighter.
2 Background and Motivation
(a) Patch for the null-pointer-dereference bug. The pointer returned by devm_kzalloc should be checked.
(b) A new bug detected by KNighter with CVE-2024-50103.
(c) The checker synthesized by KNighter.
Figure 1: A bug pattern related to devm_kzalloc.
2.1 Clang Static Analyzer
Static analysis [4] inspects code without execution to uncover bugs and vulnerabilities. The Clang Static Analyzer [8] (CSA) employs path-sensitive symbolic execution, constructing an Exploded Graph where each node (an ExplodedNode) captures a ProgramPoint and an abstract ProgramState that maps expressions to symbolic values and memory locations. Its modular design uses checkers—small, event-driven components implemented as subclasses of a Checker template—that register for events (e.g., pre- and post-function calls, dead symbols, pointer escapes) and can extend the ProgramState with custom data via dedicated macros. Developing a new checker involves defining the bug pattern, implementing the appropriate callbacks, registering the checker, and integrating it into the testing framework, all while ensuring clear bug reporting through BugType and BugReport mechanisms.
Take the checker shown in Figure 1(c) as an example. In checkPostCall, the checker first verifies that the function call originates from "devm_kzalloc" using ExprHasName, and if so, it updates the program state by marking the memory region in PossibleNullPtrMap as unchecked. In checkBranchCondition, two patterns are handled: for instance, when a pointer is negated in a condition (e.g., if (!ptr)), the checker marks the region as checked by calling markRegionChecked; similar logic applies for conditions comparing the pointer directly to NULL. The checkLocation callback then looks up the region in PossibleNullPtrMap and, if it remains unchecked, issues a warning through reportUncheckedDereference. Finally, checkBind addresses pointer assignments by updating the PtrAliasMap for both the left-hand side and right-hand side regions, ensuring that any aliasing is properly tracked.
2.2 Motivating Example
We demonstrate KNighter’s effectiveness through a case study involving a Null-Pointer-Dereference vulnerability pattern. Figure 1(a) shows a historical patch addressing this pattern, where the original bug stemmed from a missing null pointer check after a devm_kzalloc call. Without this check, the system could crash if memory allocation failed and the returned null pointer was subsequently dereferenced.
Limitations of existing tools. Despite this vulnerability pattern recurring since at least 2021, with our analysis identifying at least four historical patches addressing it, no static analysis tool had been developed to systematically detect these issues. Even specialized kernel checkers like Smatch [1] fail to identify these vulnerabilities because they lack the domain-specific knowledge that devm_kzalloc may return NULL upon failure.
Our approach. KNighter extracts critical insights from the patch: unchecked return values from devm_kzalloc represent potential null-pointer dereference vulnerabilities. The synthesized checker (Figure 1(c)) tracks null-check status across execution paths while correctly handling pointer aliasing—a sophisticated static analysis capability. This checker discovered 3 new vulnerabilities in the Linux kernel. Figure 1(b) presents one such vulnerability exhibiting the same pattern where a null pointer check is missing after a devm_kzalloc call. This bug was subsequently fixed and assigned CVE-2024-50103, confirming the security significance of our approach.
Advantages over direct LLM scanning. Directly using LLMs to scan the Linux kernel would be prohibitively expensive, as devm_kzalloc alone appears over 7K times across 5.4K files. In contrast, KNighter’s static analyzers primarily consume CPU resources rather than repeated LLM invocations, making the approach both scalable and cost-effective. Moreover, since generating the checkers is mostly a one-time effort, they can naturally evolve alongside the codebase.
Technical challenges and solutions. Creating effective static analyzers with LLMs presents several challenges. First, writing robust checkers end-to-end is complex. KNighter addresses this through a multi-stage synthesis pipeline that breaks down complex tasks into manageable steps. Second, LLM hallucination can produce incorrect analyzers. KNighter mitigates this by validating synthesized checkers against historical patches, verifying they correctly distinguish between buggy and patched code. Finally, to reduce false positives, we implement a bug triage agent that identifies false alarms, enabling iterative refinement of the checkers.
Figure 2: Overview of KNighter.
3 Design
Terminology. KNighter takes a patch commit as input and outputs a corresponding CSA checker. Valid checkers correctly distinguish between buggy and patched code, flagging pre-patch code as defective while recognizing post-patch code as correct. Plausible checkers [^1] We adopt the term “plausible” from program repair literature [24, 31], where a “plausible” patch passes all test cases and potentially represents the correct fix. are valid checkers that additionally demonstrate practical utility through low false positive rates or a manageable number of reports. We provide formal definitions of these terms in § 4.
Overview. KNighter leverages agentic workflow to process patch commits for static analyzer synthesis, as illustrated in Figure 2. It operates in two phases: checker synthesis (§ 3.1) and checker refinement (§ 3.2). In the checker synthesis phase, KNighter analyzes the input patch to identify bug patterns (§ 3.1.1), synthesizes a detection plan (§ 3.1.2), and implements a checker using Clang (§ 3.1.3). If compilation errors occur, a syntax-repair agent automatically repairs them based on the error messages. This phase concludes with the generation of valid checkers (§ 3.1.4). In the subsequent checker refinement phase, these valid checkers are deployed to scan the entire codebase for potential bugs. When bug reports are generated, a bug-triage agent evaluates them for false positives, and KNighter refines the checker accordingly. If the scan produces a manageable number of reports with a low false positive rate, KNighter presents the plausible checkers and their filtered reports as potential bugs for review.
Figure 3: Patch commit message.
3.1 Checker Synthesis
Algorithm 1 presents the multi-stage pipeline of checker synthesis. In the first stage, KNighter analyzes the bug pattern shown in the patch (Line 1). Next, KNighter synthesizes the plan based on the patch and the identified bug pattern (Line 1). With the plan in hand, KNighter implements the checker using CSA (Line 1). If any compilation issues arise, a syntax-repair agent is invoked to debug and repair them (Line 1). The repair process is allowed up to maxAttempts (default is 5) attempts. If the checker compiles successfully, KNighter validates it by checking whether it can distinguish between the buggy and patched code (Line 1). Once the checker is deemed valid, it is returned for the next phase (Line 1). Otherwise, the synthesis pipeline continues iterating until reaching maxIterations. If all iterations fail, the process returns Null, indicating that a valid checker could not be synthesized (Line 1).
1
2
3Function GenChecker(patch):
4       // Iterative checker generation and evaluation
5       for i = 1 to maxIterations do
6             // Stage 1: Bug Pattern Analysis
7             pattern $\leftarrow$ AnalyzePatch(patch)
8
9            // Stage 2: Detection Plan Synthesis
10             plan $\leftarrow$ SynthesizePlan(patch, pattern)
11
12            // Stage 3: Analyzer Implementation and Repair
13             checker $\leftarrow$ Implement(patch, pattern, plan)
14             attempts $\leftarrow$ 0
15             while hasCompilationErrors(checker) and attempts < maxAttempts do
16                   checker $\leftarrow$ RepairChecker(checker)
17                   attempts $\leftarrow$ attempts + 1
18
19
20            if hasCompilationErrors(checker) then
21                   // Skip evaluation if checker still has errors
22                   Continue
23
24
25            // Stage 4: Validation
26             isValid $\leftarrow$ ValidateChecker(checker, patch)
27
28            if isValid then
29                   return checker
30
31
32      return Null
33
Algorithm 1 Synthesize checkers with patch commits.
3.1.1 Bug Pattern Analysis
The initial stage involves analyzing patch commits to identify underlying bug patterns. Patch commits typically consist of diff patches and may include developer comments describing the bug being fixed, as illustrated in Figure 3. Our goal is to extract patterns that can be translated into static analysis rules for bug detection. While bug patterns are sometimes explicitly described in commit messages, they often require deeper analysis of the code changes within the patch.
Figure 4: Prompt template for bug pattern analysis
We have developed an LLM-based agent specifically designed to perform this pattern analysis, with the prompt template shown in Figure 4. In addition to the patch, we extract the complete function code that was modified from the kernel codebase. This additional context is crucial because the patch diff alone may not capture all relevant buggy patterns, as some issues depend on the broader context of the code. By providing both the patch and the complete function code to LLMs, we enable a more comprehensive understanding of the bug being patched.
A single bug pattern from a patch can often be expressed and detected in multiple ways, varying in their scope and complexity. Consider the null pointer dereference bug shown in Figure 1(a). One could formulate a general pattern requiring checks after any function that might return a null pointer. While comprehensive, implementing such a broad checker would be challenging even for LLMs. Instead, our approach favors more targeted bug patterns that facilitate precise checker implementation while remaining practical. For the null pointer example, we narrow the scope to specifically verify that pointers returned by devm_kzalloc are checked before dereferencing—a more focused pattern that maintains effectiveness while being significantly more tractable to implement.
3.1.2 Plan Synthesis
After identifying the bug pattern, KNighter synthesizes a high-level plan for implementing the static analyzer. This planning phase is critical for two reasons: it provides structured guidance for LLMs during implementation, thereby preventing confusion or ineffective execution, and it facilitates debugging of the entire pipeline by making the LLMs’ reasoning process transparent and traceable. Our ablation study in § 5.4.2 demonstrates that having a plan leads to better performance than not having one, consistent with findings from other domain applications [27].
Figure 5: Prompt template for plan synthesis.
To synthesize the implementation plan for the checker, we have designed an LLM-based agent whose prompt template is shown in Figure 5. This agent takes the previously summarized bug pattern as input. Additionally, we maintain a curated database of utility functions for checker implementation that can be easily extended. By including the signatures and brief descriptions of these utility functions in the prompt, we enable LLMs to leverage them effectively during the planning process, simplifying the overall task.
3.1.3 Analyzer Implementation and Syntax Repair
After identifying the bug pattern and making the plan, we leverage an LLM-based agent to implement the corresponding checker, as illustrated in Figure 6. To maximize implementation accuracy, we provide the agent with comprehensive inputs: the distilled bug pattern and a structured implementation plan. We also provide a pre-defined checker template that standardizes the implementation structure and reduces potential errors. Moreover, we provide a list of utility functions that could help with the implementation.
To handle potential compilation errors in the generated checkers, we employ a dedicated debugging agent. Inspired by existing work on program repair [32], this agent automatically processes compiler error messages and applies necessary fixes, effectively addressing syntax errors that may arise from LLM hallucinations. This automated debugging pipeline ensures that the final checkers are both syntactically correct and compilable.
Figure 6: Prompt template for analyzer implementation.
3.1.4 Validation
To validate our checkers semantically and mitigate hallucinations by LLMs, we evaluate them against both the buggy (pre-patch) and patched versions of the Linux kernel. This dual-version analysis verifies that the checker can both detect original bugs and recognize their fixes. For efficiency, we restrict the evaluation to only the files modified by the patch and their dependencies, rather than scanning the entire kernel codebase. This focused approach enables rapid validation while maintaining effectiveness. We assess checker validity by comparing the number of reported issues between versions, acknowledging that while similar bugs might persist in the patched version, an effective checker should yield a meaningful difference in detection results. Detailed implementation specifics are presented in § 4.
3.2 Checker Refinement
After synthesizing a valid checker, we scan the entire Linux codebase. While effective at identifying potential issues, these checkers may be overly conservative, flagging correct code as problematic and generating false positives. To address this limitation, we develop an iterative refinement process powered by LLMs. The process evaluates each bug report’s validity and uses false positive cases to refine the checker. However, this refinement process faces two significant challenges:
  • Bug reports often contain extensive context and implementation details, making them difficult to process efficiently.
  • Debugging and refining checkers based on false positive reports requires complex analysis and careful modification.
To address these challenges, we first distill bug reports to their essential elements. We extract only the "relevant lines" identified by the LLVM report and the trace path for each potential issue—preserving critical diagnostic information while eliminating extraneous context. We then deploy a specialized LLM-based triage agent to classify each report as either a true or false positive. Note, our classification criteria focus specifically on pattern alignment: we evaluate whether each reported issue exhibits the same bug pattern as the original input patch, rather than conducting a general correctness assessment. For reports classified as false positives, a dedicated refinement agent analyzes the specific case and modifies the checker accordingly, enhancing precision while preserving the checker’s ability to detect the original vulnerability pattern.
A refined checker is accepted only if it satisfies two criteria: (1) it no longer generates warnings for the previously identified false positive cases, and (2) it maintains its validity by correctly differentiating between the original buggy and patched code versions. This criterion ensures the semantic accuracy of the refined checkers.
4 Implementation
Input commit collection. To collect patch commits, we implemented a systematic classification and selection process. First, we established 10 distinct bug categories. We then used relevant keywords to identify potentially related commits. A commit was included in our dataset only when two authors independently agreed on its categorization. For each bug type, we initially examined the first 20 commits that matched our search criteria. We continued reviewing commits beyond the initial 20 if we hadn’t yet collected 5 qualifying commits for a given category. Our goal was to gather a minimum of 5 commits per bug type whenever possible. Table 1 presents our categorization of 10 bug types and their corresponding patch commit counts.
Table 1: Distribution of patch commits across 10 bug categories and the validity status of their synthesized checkers. “NPD” denotes “Null-Pointer-Dereference” and “UBI” indicates “Use-Before-Initialization”.
Valid
Bug Type Total Invalid Direct Refined Fail
NPD 6 1 2 2 1
Integer-Overflow 7 3 1 3 0
Out-of-Bound 6 2 4 0 0
Buffer-Overflow 5 3 2 0 0
Memory-Leak 5 2 3 0 0
Use-After-Free 7 4 2 1 0
Double-Free 8 1 5 1 1
UBI 5 1 1 3 0
Concurrency 5 2 3 0 0
Misuse 7 3 3 1 0
Total 61 22 26 11 2
Few-shot examples. We prepared three end-to-end examples for LLMs’ in-context learning. These three are patch commit 3027e7b15b02 (Null-Pointer-Dereference), 3948abaa4e2b (Use-Before-Initialization), and 4575962aeed6 (Double-Free). The design and implementation of the checker for these three commits required approximately 40 person-hours.
Utility functions. While implementing checkers for the example commits, we identified several commonly needed utility functions. We implemented 9 utility functions, including getMemRegionFromExpr, to support checker development. These utility functions are designed to be simple to implement and extend.
Valid checkers. To evaluate checker validity, we verify that it can both detect the original bug and recognize its fix. We first identify buggy objects by examining the modified files in the diff patch. Next, we check out the repository to the buggy commit (immediately preceding the patch) and scan these objects to count the number of bug reports ($N_{buggy}$). We then scan the same objects after applying the patch commit to obtain the number of remaining bug reports ($N_{patched}$). A checker is considered valid if $N_{buggy}>N_{patched}$ and $N_{patched}<T_{valid}$, where $T_{valid}$ is a threshold value, which is 50 by default.
Plausible checkers. We determine plausible checkers based on their performance when analyzing the entire Linux kernel. Our approach is founded on the principle that high-quality checkers, especially those derived from historical commits, should generate a reasonable number of actionable bug reports. A checker is classified as plausible if it either: (1) produces fewer reports than a predefined threshold $T_{{plausible}}$ (default: 20), or (2) demonstrates an acceptable false positive rate in sampled warnings.
Checker refinement. We evaluate each valid checker by scanning the entire kernel codebase independently, with execution bounded by either a one-hour time limit or a maximum of 100 warnings during the refinement process. Note that these limits are only applied during the checker refinement phase; when performing actual bug detection, we run the checkers without such constraints. The refinement process begins with LLM-assisted triage of the checker’s output. Using a consistent random seed, we sample 5 warnings for LLM inspection due to cost consideration. A checker qualifies as plausible if it either generates fewer than $T_{{plausible}}=20$ total reports or exhibits at most one false positive in the evaluated sample (labeled by our triage agent). For checkers failing these criteria, we implement an iterative refinement protocol targeting the identified false positives, permitting up to three refinement iterations to improve precision.
5 Evaluation
We explore the following research questions for KNighter:
  1. RQ-1.
    Can KNighter synthesize high-quality static analyzers?
  2. RQ-2.
    Can the checkers generated by KNighter find real-world kernel bugs?
  3. RQ-3.
    How do the static analyzers synthesized by KNighter compare with human-written ones?
  4. RQ-4.
    Are all the key components in KNighter effective?
Evaluation metrics. We conduct an extensive evaluation by using the following metrics:
Checker Validity Rate. A valid checker successfully identifies the buggy pattern in the original code and confirms its absence in the patched version. This metric reflects our framework’s and LLMs’ ability to understand patch semantics and synthesize discriminative checkers.
Plausible Checker Rate. This metric measures the number of high-quality checkers synthesized, representing those that are both valid and exhibit a low false positive rate.
Bug Detection. We assess the number of real-world bugs successfully detected by the synthesized checkers.
Resource Efficiency. This metric captures the computational time and monetary costs associated with both checker synthesis and execution.
Checker Error Categories. We classify checker failures into the following categories, ordered by severity:
  • Compilation Failures: Checkers that fail during compilation due to syntax or dependency errors.
  • Runtime Errors: Checkers that compile successfully but crash during execution (e.g., "The analyzer encountered problems on source files").
  • Semantic Issues: Checkers that cannot distinguish between the buggy and patched code.
Hardware and software. All our experiments are run on a workstation with 64 cores, 256 GB RAM, and 4 Nvidia A6000 GPUs, operating on Ubuntu 20.04.5 LTS. We use O3-mini as our default LLM backend. By default, when scanning the entire codebase, we use -j32. We evaluated using Linux v6.13, and for bug finding, we examined versions from v6.9 to v6.13. The Linux configuration used is allyesconfig.
5.1 RQ1: Synthesized Checkers
(a) The bug in the Use-Before-Initialization patch.
(b) A false positive reported by KNighter.
Figure 7: Reported false positive for a UBI commit.
Synthesis. From the commits collected in Table 1, we successfully generated valid checkers for 39 commits. Furthermore, KNighter successfully generated valid checkers across diverse bug types beyond those in our few-shot examples, demonstrating the generalizability of our approach.
The complete synthesis process required 15.9 hours, consuming 8.2 million input tokens and producing 1.2 million output tokens. For commits yielding valid checkers, KNighter required an average of 2.4 attempts per checker (maximum: 8 attempts). The generated valid checkers averaged 125.7 lines of code.
Refinement. After scanning the entire kernel codebase with these 39 valid checkers, 26 of them were labeled “plausible” directly. Our refinement pipeline was applied to the remaining 13 valid checkers, successfully refining 11 of them. In total, 19 refinement steps were completed successfully. This demonstrates the effectiveness of our refinement pipeline, which successfully refined 84.6% of the valid checkers that were not initially labeled as “plausible”.
False positive rate. Of the 37 plausible checkers, 16 did not report any bugs. For the remaining checkers, we applied our bug triage agent to filter all the reports, focusing only on those labeled as “bug” since our triage agent demonstrated a low false negative rate in our evaluation (as shown in § 5.4.1). In total, we obtained 90 reports labeled as “bug”. Upon manual verification, we confirmed 61 true positives. This indicates that the combination of our plausible checkers and bug triage agent has a false positive rate of 32.2%.
Here are some example checkers with poor performance (high false positive rate). For the commit 90ca6956d383 (“ice: Fix freeing uninitialized pointers”) shown in Figure 7(a), the issue occurs when the pointer pcaps is not set to NULL, and an early return or error path is taken before it is assigned a valid allocation, causing the cleanup logic to attempt to free an uninitialized (or garbage) pointer. However, to prove such a bug pattern exists, there must be an early exit-path where the pointer is not assigned any value. Thus, for the bug report shown in Figure 7(b), although the cert pointer is not initialized, it will be directly assigned a value and there is no path where it remains unassigned. Therefore, this is not a bug. Our synthesized checker failed to consider this constraint, and our triage agent also overlooked this issue.
Invalid checkers. During the checker synthesis process (up to 10 attempts for each commit), we recorded in total 273 failed attempts across all 61 commits. These invalid checkers failed for three primary reasons: 65 encountered compilation errors, 1 exhibited runtime errors, and the remaining 207 contained semantic issues that prevented proper bug identification.
Of the 207 checkers with semantical issues, 34 incorrectly labeled both buggy and patched code as potential bugs, while the remaining 173 incorrectly labeled both versions as not containing bugs. This demonstrates that correctly identifying buggy code remains a significant challenge for KNighter.
Interestingly, the 173 checkers that fail to identify the specific bug in their input patch can still provide value. When applied to the broader kernel codebase, these checkers may successfully detect bugs with similar patterns in other contexts. This apparent paradox occurs because a checker’s failure to detect its training bug might be due to specific edge cases or context-dependent factors rather than fundamental flaws in its detection logic. For instance, a checker might not detect a null pointer dereference in its training example due to complex control flow, but still correctly identify simpler instances of the same pattern elsewhere. Moreover, these checkers often have lower false positive rates compared to those that incorrectly flag both buggy and patched code as problematic, making them even more practical for real-world bug detection.
Our approach shows limitations in detecting buffer overflow and use-after-free vulnerabilities. A key challenge is that CSA has inherent difficulty in performing precise value analysis, particularly in determining buffer bounds at compile time. This limitation makes it challenging to statically reason about potential overflow conditions that depend on runtime values.
Table 2: Detected bugs by KNighter.
Total Confirmed Fixed Pending CVE
KNighter 70 56 41 15 11
(a) Number of bugs in each subsystem.
(b) Number of bugs in each type.
Figure 8: The details of newly detected bugs by KNighter.
(a) Input Use-After-Free patch.
(b) CVE-2025-21715, found by the checker for the patch above.
(c) Input Buffer-Overflow patch.
(d) CVE-2024-50259, found by the checker for the patch above.
Figure 9: Example vulnerabilities detected by KNighter.
5.2 RQ2: Detected Bugs
To date, the static analyzers generated by KNighter have detected 70 new bugs, with 56 confirmed and 41 already fixed (see Table 2). Notably, 11 of these bugs have been assigned CVE numbers. We identified more bugs than those reported as true positives in RQ-1 because our analysis spanned a broader range of Linux kernel versions.
Figure 8 shows the details of newly detected bugs, including their subsystem locations and bug types. We found that most bugs are from the drivers subsystem (45/70), which aligns with the fact that driver code constitutes the largest portion of the kernel [6]. Following drivers, we also found 12 and 6 bugs in the sound and net subsystems respectively.
As shown in Figure 8(b), KNighter demonstrates capability in detecting a diverse range of bug types. Null-Pointer-Dereference issues constituted the largest category (30/70 bugs), followed by Integer-Overflow and Misuse vulnerabilities. This distribution likely reflects the pattern-based nature of Null-Pointer-Dereference and Misuse bugs, which makes their checkers more amenable to LLM-based implementation. Integer-Overflow bugs, typically involving variable type mismatches, benefit from the Clang toolchain’s strong type analysis capabilities. More complex issues such as Use-After-Free and Concurrency bugs proved more challenging to detect at compile time, particularly those involving multi-threaded execution contexts. Overall, KNighter-synthesized checkers are general to detect various types of bugs.
Here are example vulnerabilities detected by KNighter.
CVE-2025-21715. Figure 9(a) (the input patch to KNighter) shows a fix for a Use-After-Free vulnerability. In this patch, free_netdev must be invoked only after all the references to its private data, otherwise, it could cause a Use-After-Free issue. Leveraging this patch, the checker generated by KNighter identified a similar bug in dm9000_drv_remove, as shown in Figure 9(b), where dm (the private data of ndev) remains in use after ndev is freed, causing a Use-After-Free. This newly discovered issue was assigned CVE-2025-21715.
CVE-2024-50259. Figure 9(c) shows an input patch addressing a buffer overflow vulnerability. The patch mitigates the risk by limiting the number of bytes copied via copy_from_user to sizeof(mybuf) - 1, thereby preserving space for a trailing zero. This trailing zero is essential for subsequent string operations, such as sscanf, to function correctly. Leveraging this patch as a reference, the checker generated by KNighter identified a similar bug in nsim_nexthop_bucket_activity_write, as shown in Figure 9(d). In this case, the omission of appending a trailing zero after copying data from userspace could lead to improper string handling and potential overflow issues. This detected issue was subsequently fixed by adding the trailing zero and was assigned CVE-2024-50259.
5.3 RQ3: Baseline Comparison
Since no comparable automated static analyzer generation approaches exist for this domain, we evaluate KNighter against expert-written checkers. Our baseline is provided by Smatch [1], which is widely used in Linux kernel analysis and supports tailored checks for all bug types considered in Table 1. We conducted the comparative analyses by running Smatch on the entire codebase to determine if it could detect the bugs found by KNighter.
Our evaluation revealed that Smatch reported 1970 errors and 2870 warnings across the kernel. Notably, it failed to detect any of our true positive bugs, underscoring KNighter’s unique detection capabilities. Further investigation of Smatch’s checkers showed they don’t fully leverage domain-specific knowledge available in the Linux kernel—knowledge that KNighter extracts from historical patches. For instance, Smatch’s check_deref checker employs static range analysis to identify potential null pointers but lacks domain-specific insights. It fails to recognize that functions like devm_kzalloc may return NULL under error conditions that conventional static range analysis cannot detect. Consequently, Smatch identified only three potential null pointer dereferences, all of which were confined to unit test files rarely prioritized by developers. We conclude that KNighter and Smatch detect different classes of bugs, demonstrating KNighter’s effectiveness in learning domain knowledge from patches and subsequently identifying diverse bugs and vulnerabilities.
5.4 RQ4: Effectiveness of Components
5.4.1 Bug Triage Agent
To evaluate our triage agent, from the valid checkers, we sampled up to 5 reports per checker to reduce manual inspection efforts. In total, we collected 79 reports from 18 checkers. The remaining 21 valid checkers did not identify any bugs, while 2 non-plausible checkers also generated reports. These reports may include those generated during the refinement process. Our objective here was to evaluate the effectiveness of our triage agent. Among these reports, 29 were labeled as “bug” (positive) by our triage agent, while the remaining 50 were labeled as “not-a-bug” (negative). Two of our authors manually cross-checked all reports to establish ground truth. Our triage agent achieved 7 true positives (TP), 22 false positives (FP), 50 true negatives (TN), and zero false negatives (FN). Notably, the absence of false negatives is beneficial because it helps ensure that most true bugs are captured, even if this comes with some false positives that can be filtered out in subsequent steps or human inspection.
5.4.2 Ablation Study
Table 3: Ablation study results. “Default” means KNighter’s standard configuration utilizing multi-stage synthesis and the O3-mini language model. Alternative configurations are compared against this baseline.
Variants Valid Errors
Syntax Runtime Semantics
Default 12 28 0 75
W/o multi-stage 8 52 3 75
W/ GPT-4o 11 31 0 76
W/ Gemini-2-flash 4 130 2 44
To further evaluate our design choices, we created a sample dataset of patch commits for an ablation study. We randomly sampled 2 commits from each bug type using zero as the random seed. This resulted in a dataset of 20 commits (2 commits × 10 bug types). Table 3 shows the results of our ablation study.
Checker synthesis. Our default pipeline implements a three-stage approach: bug pattern analysis, detection plan synthesis, and checker implementation. We evaluated the effectiveness of directly synthesizing checkers without the initial bug pattern analysis and plan synthesis phases while maintaining identical in-context learning examples across configurations. As shown in Table 3, without multi-stage synthesis, KNighter produced valid checkers for only 8 commits compared to 12 with our default approach. The single-stage approach also generated significantly more syntax errors (52 vs. 28), resulting in checkers that failed to compile.
LLM choice. We evaluated KNighter’s performance across different language models for checker synthesis. In addition to our default model O3-mini (reasoning model), we tested GPT-4o and Gemini-2-flash. As Table 3 demonstrates, GPT-4o performed slightly worse than O3-mini, generating 11 valid checkers compared to 12. Gemini-2-flash performed substantially worse, producing valid checkers for only 4 commits. Upon closer inspection, we found that Gemini-2-flash struggled with Clang Static Analyzer implementation, frequently using non-existent APIs and generating syntax errors at a much higher rate (130 vs. 28).
6 Related Work
6.1 Traditional Static Analysis
Given the cruciality of Linux kernel and the diversity of its vulnerabilities, many static analyzers have been developed to target different classes of bugs.
CRIX [20] focuses on detecting missing-check bugs in kernel by constructing def-use slices of critical variables. Goshawk [21] proposes a structure-aware and object-centric abstraction over memory management operations to aid detecting double-free and use-after-free bugs. UBITect [39] integrates type qualifier analysis and symbolic execution to detect use-before-initialization bugs in the Linux kernel, which might encounter challenges due to exhaustive path exploration. CRed [34] detects Use-After-Free bugs in scale by spatio-temporal context reduction. LR-Miner [16] establishes field-sensitive locking rules between locks and variables. Then it detects rule violations to pinpoint data races in mainstream OS kernels. Authors of [12] extensively analyze over one thousand reference counting bugs in the history of Linux kernel and derived 9 anti-patterns. They further developed a set of static analyzers to detect such patterns. LinkRID [19] also targets refcounting bugs by modelling common Linux-specific refcount usage conventions. DCUAF[5] conducts local-global analysis to find concurrent functions to discover kernel currency bugs. SUTURE[41] performs high-order taint analysis to discover kernel vulnerabilities caused by malicious userspace input.
While traditionally designed static analyzers can perform well on their targeted classes of bugs, they require extensive analysis and implementation effort from human experts. This limits their ability to scale to additional bug classes and keep pace with the rapid evolution of the Linux kernel. KNighter is complementary to such analyzers by detecting more classes of bugs and adapting to newly developed bug patterns.
Beyond one specific class of bugs, many static analysis systems implement different analysis techniques to improve precision and efficiency. FiTx [28] implements fast analysis for single-compilation unit without the requirement of complex features like indirect call analysis. PATA [17] improves the alias analysis precision using control-flow paths and access paths information. FiTx and PATA support detecting 6 and 3 common bug patterns respectively, by encoding bug patterns as finite state machines. While these static analysis systems can potentially support more classes of bugs, analyzing bugs and designing the pattern state machines remain challenging and labor-intensive. KNighter automates this process by automatically synthesizing bug patterns from patches. Beyond automation, although KNighter generates CSA checkers, the idea is generalizable to using other static analysis tools without fundamental obstacles since LLMs have been trained on most existing tools.
6.2 LLM-Based Static Analysis
With the recent advances of LLMs, many techniques leverage LLMs’ code comprehension ability to mediate the limitations of traditional static analysis tools that typically require approximation or expert-produced knowledge. IRIS [18] enhanced CodeQL’s taint analysis by using LLMs to infer taint specifications for external APIs, which would have required manual labeling. To detect Use-Before-Initialization bugs, LLift [15] uses LLMs to extract initializers and generate their post-constraints. The post-constraints are then used during further analysis to prune unreachable paths. InferROI [30] uses LLMs to infer resource-oriented intentions and then perform lightweight static resource leak detection. While the above approaches can enhance static analyzers, a substantial portion of the analyzer still needs to be crafted by researchers. In contrast, KNighter can automatically synthesize the entire checker without human intervention. Beyond this, the checkers generated by KNighter could potentially be further complemented by additional information provided by the above approaches.
Another branch of work directly use LLMs to analyze the code without traditional static analysis tools. Vul-RAG [10] constructs a knowledge base from existing CVE instances and directly queries LLMs with Retrieval-Augmented-Generation to determine whether a new piece of code has the same bug. Zhang et al. [40] studies various prompt design techniques for vulnerability detection. Fine-tuning has also been explored by [25] to perform vulnerability detection. The above LLM-Based analysis techniques require querying LLMs with the code under test. In the context of Linux kernel, querying LLMs with millions lines of code, potentially multiplied by the number of different classes of bugs, is prohibitively expensive. In comparison, KNighter generates a static analysis checker with negligible token cost that can be used to efficiently scan the kernel codebase.
7 Conclusion
This paper introduces KNighter, a novel approach that transforms how LLMs can contribute to static analysis for complex systems like the Linux kernel. By synthesizing specialized static analyzers rather than directly analyzing code, KNighter bridges the gap between LLMs’ reasoning capabilities and the practical constraints of analyzing massive codebases. KNighter’s practical impact is demonstrated by the discovery of 70 previously unknown bugs in the Linux kernel, with 56 confirmed and 41 already fixed.
Looking forward, KNighter opens new possibilities for practical LLM-based static analysis. Future work could extend this approach to other systems beyond the Linux kernel, incorporate additional learning paradigms, and further refine checker generation techniques to address more complex bug patterns. By leveraging LLMs to synthesize tools rather than perform analysis directly, we establish a scalable, reliable, and traceable paradigm for utilizing AI in critical software security applications.

### v2

@@
 KNighter: Transforming Static Analysis with LLM-Synthesized Checkers
-Abstract
+Abstract.
 Static analysis is a powerful technique for bug detection in critical systems like operating system kernels.
 However, designing and implementing static analyzers is challenging, time-consuming, and typically limited to predefined bug patterns.
-While large language models (LLMs) have shown promise for static analysis, directly applying them to scan large codebases remains impractical due to computational constraints and contextual limitations.
-We present KNighter, the first approach that unlocks practical LLM-based static analysis by automatically synthesizing static analyzers from historical bug patterns.
-Rather than using LLMs to directly analyze massive codebases, our key insight is leveraging LLMs to generate specialized static analyzers guided by historical patch knowledge.
+While large language models (LLMs) have shown promise for static analysis, directly applying them to scan large systems remains impractical due to computational constraints and contextual limitations.
+We present KNighter, the first approach that unlocks scalable LLM-based static analysis by automatically synthesizing static analyzers from historical bug patterns.
+Rather than using LLMs to directly analyze massive systems, our key insight is leveraging LLMs to generate specialized static analyzers guided by historical patch knowledge.
 KNighter implements this vision through a multi-stage synthesis pipeline that validates checker correctness against original patches and employs an automated refinement process to iteratively reduce false positives.
 Our evaluation on the Linux kernel demonstrates that KNighter generates high-precision checkers capable of detecting diverse bug patterns overlooked by existing human-written analyzers.
-To date, KNighter-synthesized checkers have discovered 70 new bugs/vulnerabilities in the Linux kernel, with 56 confirmed and 41 already fixed.
-11 of these findings have been assigned CVE numbers.
+To date, KNighter-synthesized checkers have discovered 92 new, critical, long-latent bugs (average 4.3 years) in the Linux kernel; 77 are confirmed, 57 fixed, and 16 have been assigned CVE numbers.
 This work establishes an entirely new paradigm for scalable, reliable, and traceable LLM-based static analysis for real-world systems via checker synthesis.
-1 Introduction
-The reliability of fundamental systems, particularly operating system kernels, depends critically on effective defect detection techniques [23, 38, 12, 7, 21, 28, 35].
-Static analysis [4] stands out as a powerful approach that can comprehensively examine source code without execution, enabling the detection of bugs in architecture-specific code, hardware-dependent drivers, and configurations that may be challenging to test in real environments [20, 21, 12, 28].
-Unlike dynamic approaches such as fuzzing that require concrete execution environments and can only explore executed paths [2, 38, 9, 36], static analysis can theoretically examine all possible code paths, including those that are difficult or impossible to trigger during runtime.
-Plus, while formal verification [37, 14, 26] offers mathematical guarantees that static analysis cannot provide, it typically requires even more expertise and manual effort, making it impractical for large-scale systems like operating system kernels.
-However, static analysis faces its own set of challenges in practice.
-Developing effective static analyzers requires deep expertise in both the target system and formal methods, along with substantial engineering effort to implement precise detection mechanisms [12, 7, 1].
-This complexity often leads to static analyzers that focus on specific classes of bugs, potentially missing other critical defect patterns.
-These observations highlight several key opportunities for advancing static analysis.
-First, the rich history of the fundamental systems, e.g., the Linux kernel, provides extensive examples of bug patterns that can inform static analysis approaches.
-Moreover, bug patterns identified in historical patch commits may still exist in the current codebase [12].
-Third, beyond manually analyzing and summarizing bug patterns, modern large language models (LLMs) [3, 29, 11] have demonstrated remarkable capabilities in understanding both code and natural language descriptions.
-Building on these insights, we propose leveraging LLMs to perform static analysis on real-world systems by learning from historical patch commits.
-However, directly applying LLMs to analyze real-world systems [38, 15, 18, 42, 10, 33] is challenging due to the enormous size of modern codebases—for example, the Linux kernel comprises over 30 million lines of code.
-The limited context windows of LLMs make it impractical to process an entire codebase at once, and the associated computational cost is prohibitive.
-In the case of the Linux kernel, which spans more than 300 million tokens, a single analysis pass using GPT-4o is estimated to cost approximately \$750.
-Moreover, LLMs are susceptible to hallucination issues, potentially leading to false positives [13, 10, 18].
-Besides, LLMs also struggle to understand complex codebases [22].
-Our solution addresses these challenges by using LLMs to synthesize static checkers based on historical bug patches rather than directly analyzing source code.
-These checkers, implemented using Clang Static Analyzer (CSA) [8], can then scan the kernel codebase effectively and efficiently.
-This approach offers several advantages:
-  • It avoids the prohibitive costs and context limitations of processing large codebases directly with LLMs, thereby enhancing scalability.
-  • The checkers can naturally evolve alongside kernel development without requiring constant LLM involvement.
-  • We can validate the synthesized checkers against the original patches to mitigate hallucination concerns.
-  • Our approach provides greater traceability and explainability compared to earlier techniques, as the synthesized checkers directly encode the bug patterns in a transparent, human-readable manner.
-However, we observed that current LLMs struggle to synthesize and implement checkers end-to-end—a task that challenges even experienced researchers and developers with expertise in both static analysis and the target codebase.
-To address this limitation, we develop a multi-stage synthesis pipeline (§ 3.1) that breaks down the complex task into manageable steps.
-Furthermore, we design a fully automated refinement pipeline (§ 3.2) that iteratively improves the synthesized checkers to reduce false positives, ensuring their practical utility in real-world scenarios.
-We implement our approach, KNighter, the first fully automated static analyzer generation pipeline.
-While our approach is highly generalizable to different software systems, we demonstrate its effectiveness by synthesizing checkers for 61 historical patch commits from the Linux kernel, successfully generating 37 plausible checkers.
-These checkers demonstrate high quality, with approximately 35% false positive rate.
-Our evaluation also presents that KNighter-detected vulnerabilities are orthogonal to existing expert-written static analyzers [1].
-To date, checkers synthesized by KNighter have detected 70 new bugs/vulnerabilities in the Linux kernel, with 56 confirmed and 41 already fixed.
-Notably, 11 of these bugs have been assigned CVE numbers.
-Our contributions are summarized below:
-  • New paradigm.
-We introduce a pioneering LLM-based approach for synthesizing static analyzers from patch commits.
-To our knowledge, KNighter is the first fully automated static analyzer generation system, establishing a new paradigm for LLM-powered static analysis.
-  • End-to-end pipeline.
-We implement KNighter as a comprehensive system featuring multi-stage synthesis and automated refinement pipelines for the Linux kernel.
-This design enables detection of diverse bug classes and adapts dynamically to emerging vulnerability patterns.
-  • Systematic evaluation.
-We rigorously evaluate KNighter by synthesizing checkers for the patch commits spanning multiple bug categories.
+1. Introduction
+Figure 1.
+Motivation.
+Static analysis should scale to address a diverse range of bug patterns and handle massive codebases.
+Traditional static analysis struggles with covering wide-ranging bugs, whereas LLM-based methods face hurdles in scaling to large codebases.
+The reliability of fundamental software systems—particularly operating system (OS) kernels—hinges on robust defect detection methodologies (kernel-fuzzing, ; kernelgpt, ; kernel-static-refcounting, ; kernel-static-cai2023place, ; kernel-static-lyu2022goshawk, ; suzuki2024fitx, ; yang2024whitefox, ; gong2025snowplow, ; gong2023snowcat, ).
+Among various techniques, static analysis (ayewah2008using, ) stands out for its ability to examine source code without execution, making it indispensable for scenarios involving hardware-dependent drivers, complex or rarely exercised paths, and configurations difficult to reproduce in real environments (kernel-static-lu2019detecting, ; kernel-static-lyu2022goshawk, ; kernel-static-refcounting, ; suzuki2024fitx, ; gong2025snowplow, ; gong2023snowcat, ).
+Compared to dynamic approaches such as fuzzing—which requires an execution environment and thus only tests actual runtime paths (syzkaller, ; kernelgpt, ; deng2023large, ; yang2023fuzzing, )—static analysis can (in principle) cover all potential execution paths, including corner cases that are seldom triggered in practice.
+While formal verification techniques (yang2024autoverus, ; lattuada2023verus, ; sun2024clover, ) offer stronger correctness guarantees, their high manual overhead renders them impractical for large-scale systems like OS kernels, making static analysis a more scalable and feasible solution.
+The static analysis problem.
+Large-scale systems present a dual challenge for static analysis: addressing diverse bug patterns and managing enormous codebases, as illustrated in Figure 1.
+An ideal static analyzer should (i) detect a wide range of defects—including those related to nuanced, system-specific semantics—and (ii) efficiently process millions of lines of code.
+However, existing techniques typically compromise on one of these critical objectives.
+Traditional static analysis.
+Traditional static analyzers are effective in identifying certain bug types, yet they fundamentally rely on pre-defined, rule-based, or formally modeled checks.
+This reliance necessitates extensive domain expertise and substantial engineering effort for their development and maintenance (kernel-static-refcounting, ; kernel-static-cai2023place, ; smatch, ).
+Consequently, these tools are often fine-tuned to a narrow subset of bug patterns, which not only limits their ability to detect unforeseen defects but also hampers their scalability in automatically addressing a broader spectrum of issues.
+Emerging LLM-based static analysis.
+On the other hand, Large Language Models (LLMs) are compelling tools for discovering bug patterns in part because they can learn directly from historical patch commits—a treasure trove of real fixes and associated bug contexts (kernel-static-refcounting, ; seal, ; aphp, ).
+Their ability to parse both textual and code content (achiam2023gpt, ; team2023gemini, ; grattafiori2024llama, ) suggests that LLMs can adapt to new bug types without explicit rule-crafting.
+However, directly deploying LLMs on large-scale systems (e.g., the Linux kernel at over 30 million lines of code) confronts severe limitations.
+Their bounded context windows make it impossible to upload all relevant source code at once, and doing so repeatedly would also incur prohibitive computational costs (potentially hundreds of dollars per thorough scan).
+In addition, LLMs can hallucinate (huang2025survey, ; du2024vul, ; li2024llm, ), producing plausible but incorrect outputs, especially when faced with the intricacies of large-scale systems (mathai2024kgym, ).
+Insight.
+Can we scale and automate static analysis to handle both diverse bug patterns and enormous codebases?
+We answer this question by harnessing the strengths of traditional static analysis alongside emerging LLMs.
+More specifically, we propose synthesizing static checkers using LLMs rather than applying LLM-based analysis directly to the entire codebase.
+In this paradigm, LLMs learn bug patterns from historical patches, and these insights are encoded into dedicated static analysis checkers.
+This method circumvents the prohibitive costs and context-length limitations of scanning vast codebases while maintaining the flexibility needed to address a wide spectrum of bugs.
+Moreover, by validating each synthesized checker against the original patches, we mitigate hallucinations and produce transparent, human-readable logic that developers can trust and maintain.
+Technical challenges and our solutions.
+Although automated checker synthesis holds significant promise, generating complete static analysis logic remains a formidable challenge—even experts struggle with it.
+To address this, we introduce a multi-stage synthesis pipeline (§ 3.1) that decomposes checker generation into manageable subtasks.
+Furthermore, to enhance the quality of the synthesized checkers by reducing false positives, we develop a fully automated refinement pipeline (§ 3.2) that leverages bug report triage agents.
+Together, these pipelines yield checkers that are robust and practical for deployment in real-world scenarios.
+We implement our approach in a tool, KNighter, the first fully automated pipeline for synthesizing static analyzers, built upon the open-source Clang Static Analyzer (CSA) (clang-static-analysis, ).
+Although the methodology generalizes to different systems, our evaluation targets the Linux kernel, one of the most fundamental software systems.
+From 61 historical patch commits, KNighter successfully synthesizes 37 high-quality checkers with a false positive rate of about 35%.
+Furthermore, the vulnerabilities detected are orthogonal to those found by existing expert-written analyzers (smatch, ).
+Demonstrating practical impact, our synthesized checkers have uncovered 92 new, long-latent vulnerabilities (average 4.3 years) in the Linux kernel, resulting in 77 developer confirmations, 57 fixes, and 16 CVEs.
+These findings validate our approach’s efficacy and its contribution to system reliability.
+Our main contributions are summarized as follows:
+  • Novelty.
+We introduce a pioneering approach for synthesizing static analyzers from patch commits.
+To our knowledge, KNighter is the first fully automated static analyzer generation system, establishing a new paradigm for LLM-based static analysis.
+  • Approach. We implement KNighter with multi-stage synthesis and automated refinement pipelines for the Linux kernel. This design enables detection of diverse bug classes in large-scale systems.
+  • Evaluation.
+We evaluate KNighter by synthesizing checkers for the patch commits spanning multiple bug categories.
 Our results demonstrate that KNighter generates high-precision checkers with practical false positive rates.
   • Real-world impact.
-KNighter-generated checkers have discovered 70 previously unknown bugs/vulnerabilities in the Linux kernel, with 56 confirmed by developers, 41 already fixed, and 11 assigned CVE numbers—validating both the approach’s effectiveness and practical security value.
-Our code is available at ise-uiuc/KNighter.
-2 Background and Motivation
-(a) Patch for the null-pointer-dereference bug. The pointer returned by devm_kzalloc should be checked.
+KNighter-generated checkers have discovered 92 new, long-latent (average 4.3 years) bugs in the Linux kernel, with 77 confirmed, 57 fixed, and 16 assigned CVE numbers—demonstrating its practical impact on system reliability and security.
+KNighter is available at ise-uiuc/KNighter.
+2. Background and Motivation
+(a) Patch for a Null-Pointer-Dereference bug. The pointer returned by devm_kzalloc should be checked.
 (b) A new bug detected by KNighter with CVE-2024-50103.
-(c) The checker synthesized by KNighter.
-Figure 1: A bug pattern related to devm_kzalloc.
-2.1 Clang Static Analyzer
-Static analysis [4] inspects code without execution to uncover bugs and vulnerabilities.
-The Clang Static Analyzer [8] (CSA) employs path-sensitive symbolic execution, constructing an Exploded Graph where each node (an ExplodedNode) captures a ProgramPoint and an abstract ProgramState that maps expressions to symbolic values and memory locations.
-Its modular design uses checkers—small, event-driven components implemented as subclasses of a Checker template—that register for events (e.g., pre- and post-function calls, dead symbols, pointer escapes) and can extend the ProgramState with custom data via dedicated macros.
-Developing a new checker involves defining the bug pattern, implementing the appropriate callbacks, registering the checker, and integrating it into the testing framework, all while ensuring clear bug reporting through BugType and BugReport mechanisms.
-Take the checker shown in Figure 1(c) as an example.
-In checkPostCall, the checker first verifies that the function call originates from "devm_kzalloc" using ExprHasName, and if so, it updates the program state by marking the memory region in PossibleNullPtrMap as unchecked.
-In checkBranchCondition, two patterns are handled: for instance, when a pointer is negated in a condition (e.g., if (!ptr)), the checker marks the region as checked by calling markRegionChecked; similar logic applies for conditions comparing the pointer directly to NULL.
-The checkLocation callback then looks up the region in PossibleNullPtrMap and, if it remains unchecked, issues a warning through reportUncheckedDereference.
-Finally, checkBind addresses pointer assignments by updating the PtrAliasMap for both the left-hand side and right-hand side regions, ensuring that any aliasing is properly tracked.
-2.2 Motivating Example
+(c) A checker synthesized by KNighter for the patch in Fig. 2(a).
+Figure 2. A bug pattern related to devm_kzalloc.
+2.1. Clang Static Analyzer
+Static analysis (ayewah2008using, ) is a technique for detecting bugs by inspecting code without executing it.
+The Clang Static Analyzer (clang-static-analysis, ) (CSA) serves as a powerful engine for this purpose.
+It operates using path-sensitive symbolic execution, building an internal representation called an ExplodedGraph.
+Each node within this graph, an ExplodedNode, represents a specific ProgramPoint paired with an abstract ProgramState.
+This state meticulously maps program expressions to symbolic values and tracks the contents of memory locations.
+The modularity of CSA is built upon checkers.
+These are small, specialized components, typically implemented as subclasses of a Checker template.
+Checkers function in an event-driven manner, registering interest in specific analysis events such as pre- and post-function calls, the identification of dead symbols, or instances of pointer escapes.
+A key capability of checkers is their ability to extend the ProgramState with custom, checker-specific data using provided macros, allowing them to maintain sophisticated state across the analysis.
+Developing a new checker for CSA generally involves several steps: defining the specific bug pattern to be detected, implementing callback methods corresponding to the relevant analysis events, registering the new checker with the analysis framework, and integrating it into the testing system.
+Effective bug reporting is crucial, utilizing mechanisms like BugType and BugReport to provide clear diagnostics.
+To illustrate, consider the example checker shown in Figure 2(c), which registers four distinct callback functions.
+  • The checkPostCall callback activates after function calls.
+It uses ExprHasName to check if the call was to devm_kzalloc.
+If so, it updates the custom state map PossibleNullPtrMap to mark the returned memory region as potentially null (unchecked).
+  • The checkBranchCondition callback is used to handle conditional checks involving the pointer.
+It recognizes patterns like negation (if (!ptr)) or direct comparison (if (ptr == NULL)) and updates the state via markRegionChecked to reflect that a null check has occurred for the associated memory region.
+  • The checkLocation callback is triggered when a memory location is accessed.
+More specifically, it consults the PossibleNullPtrMap state; if the region is marked as unchecked at this point, it issues a warning using reportUncheckedDereference.
+  • The checkBind callback manages pointer assignments.
+It updates another custom state map, PtrAliasMap, to track potential aliases between memory regions involved in the assignment, ensuring the checker correctly handles cases where multiple pointers might refer to the same potentially null memory.
+2.2. Motivating Example
 We demonstrate KNighter’s effectiveness through a case study involving a Null-Pointer-Dereference vulnerability pattern.
@@
 Limitations of existing tools.
-Despite this vulnerability pattern recurring since at least 2021, with our analysis identifying at least four historical patches addressing it, no static analysis tool had been developed to systematically detect these issues.
-Even specialized kernel checkers like Smatch [1] fail to identify these vulnerabilities because they lack the domain-specific knowledge that devm_kzalloc may return NULL upon failure.
+Despite this vulnerability pattern recurring since at least 2017 (commit 49af64e), with our analysis identifying at least six historical patches addressing it, no static analysis tool had been developed to systematically detect these issues.
+Even specialized kernel checkers like Smatch (smatch, ) fail to identify these vulnerabilities because they lack the domain-specific knowledge that devm_kzalloc may return NULL upon failure.
 Our approach.
-KNighter extracts critical insights from the patch: unchecked return values from devm_kzalloc represent potential null-pointer dereference vulnerabilities.
-The synthesized checker (Figure 1(c)) tracks null-check status across execution paths while correctly handling pointer aliasing—a sophisticated static analysis capability.
+KNighter extracts critical insights from the patch: unchecked return values from devm_kzalloc represent potential Null-Pointer-Dereference vulnerabilities.
+The synthesized checker (written in CSA, Figure 2(c)) tracks null-check status across execution paths while correctly handling pointer aliasing, a sophisticated static analysis capability.
 This checker discovered 3 new vulnerabilities in the Linux kernel.
-Figure 1(b) presents one such vulnerability exhibiting the same pattern where a null pointer check is missing after a devm_kzalloc call.
-This bug was subsequently fixed and assigned CVE-2024-50103, confirming the security significance of our approach.
+Figure 2(b) presents one such vulnerability exhibiting the same pattern where a null pointer check is missing for the pointer returned by the devm_kzalloc call.
+This bug was subsequently fixed and assigned CVE-2024-50103.
 Advantages over direct LLM scanning.
@@
 In contrast, KNighter’s static analyzers primarily consume CPU resources rather than repeated LLM invocations, making the approach both scalable and cost-effective.
-Moreover, since generating the checkers is mostly a one-time effort, they can naturally evolve alongside the codebase.
+Moreover, since generating the checkers is mostly a one-time effort, they can naturally evolve alongside the system.
 Technical challenges and solutions.
@@
 Finally, to reduce false positives, we implement a bug triage agent that identifies false alarms, enabling iterative refinement of the checkers.
-Figure 2: Overview of KNighter.
-3 Design
+Figure 3. Overview of KNighter.
+3. Design
 Terminology.
@@
 Valid checkers correctly distinguish between buggy and patched code, flagging pre-patch code as defective while recognizing post-patch code as correct.
-Plausible checkers [^1] We adopt the term “plausible” from program repair literature [24, 31], where a “plausible” patch passes all test cases and potentially represents the correct fix.
+Plausible checkers [^1] We adopt the term “plausible” from program repair (qi2015analysis, ; xia2023automated, ), where a “plausible” patch passes all test cases and potentially is the correct fix.
 are valid checkers that additionally demonstrate practical utility through low false positive rates or a manageable number of reports.
@@
 It operates in two phases: checker synthesis (§ 3.1) and checker refinement (§ 3.2).
-In the checker synthesis phase, KNighter analyzes the input patch to identify bug patterns (§ 3.1.1), synthesizes a detection plan (§ 3.1.2), and implements a checker using Clang (§ 3.1.3).
+In the checker synthesis phase, KNighter analyzes the input patch to identify bug patterns (§ 3.1.1), synthesizes a detection plan (§ 3.1.2), and implements a checker using CSA (§ 3.1.3).
 If compilation errors occur, a syntax-repair agent automatically repairs them based on the error messages.
@@
 When bug reports are generated, a bug-triage agent evaluates them for false positives, and KNighter refines the checker accordingly.
-If the scan produces a manageable number of reports with a low false positive rate, KNighter presents the plausible checkers and their filtered reports as potential bugs for review.
-Figure 3: Patch commit message.
-3.1 Checker Synthesis
+If the scan produces a manageable number of reports with a low false positive rate, KNighter presents the plausible checkers and the filtered reports as potential bugs for review.
+Figure 4. Patch commit message.
+3.1. Checker Synthesis
 Algorithm 1 presents the multi-stage pipeline of checker synthesis.
@@
 3Function GenChecker(patch):
-4       // Iterative checker generation and evaluation
+4       # Iterative checker generation and evaluation
 5       for i = 1 to maxIterations do
-6             // Stage 1: Bug Pattern Analysis
+6             # Stage 1: Bug Pattern Analysis
 7             pattern $\leftarrow$ AnalyzePatch(patch)
 8
-9            // Stage 2: Detection Plan Synthesis
+9             # Stage 2: Detection Plan Synthesis
 10             plan $\leftarrow$ SynthesizePlan(patch, pattern)
 11
-12            // Stage 3: Analyzer Implementation and Repair
+12             # Stage 3: Analyzer Implementation and Repair
 13             checker $\leftarrow$ Implement(patch, pattern, plan)
 14             attempts $\leftarrow$ 0
-15             while hasCompilationErrors(checker) and attempts < maxAttempts do
+15             while hasCompilationErrors(checker) and attempts ¡ maxAttempts do
 16                   checker $\leftarrow$ RepairChecker(checker)
@@
 20            if hasCompilationErrors(checker) then
-21                   // Skip evaluation if checker still has errors
+21                   # Skip evaluation if checker still has errors
 22                   Continue
@@
 24
-25            // Stage 4: Validation
+25             # Stage 4: Validation
 26             isValid $\leftarrow$ ValidateChecker(checker, patch)
@@
 33
-Algorithm 1 Synthesize checkers with patch commits.
-3.1.1 Bug Pattern Analysis
+Algorithm 1 Synthesize checkers with input patch.
+3.1.1. Bug Pattern Analysis
 The initial stage involves analyzing patch commits to identify underlying bug patterns.
@@
 While bug patterns are sometimes explicitly described in commit messages, they often require deeper analysis of the code changes within the patch.
-Figure 4: Prompt template for bug pattern analysis
-We have developed an LLM-based agent specifically designed to perform this pattern analysis, with the prompt template shown in Figure 4.
+We have developed an LLM-based agent specifically designed to perform this pattern analysis, with the prompt template shown in Figure 5(a).
 In addition to the patch, we extract the complete function code that was modified from the kernel codebase.
@@
 By providing both the patch and the complete function code to LLMs, we enable a more comprehensive understanding of the bug being patched.
-A single bug pattern from a patch can often be expressed and detected in multiple ways, varying in their scope and complexity.
-Consider the null pointer dereference bug shown in Figure 1(a).
-One could formulate a general pattern requiring checks after any function that might return a null pointer.
-While comprehensive, implementing such a broad checker would be challenging even for LLMs.
-Instead, our approach favors more targeted bug patterns that facilitate precise checker implementation while remaining practical.
-For the null pointer example, we narrow the scope to specifically verify that pointers returned by devm_kzalloc are checked before dereferencing—a more focused pattern that maintains effectiveness while being significantly more tractable to implement.
-3.1.2 Plan Synthesis
-After identifying the bug pattern, KNighter synthesizes a high-level plan for implementing the static analyzer.
-This planning phase is critical for two reasons: it provides structured guidance for LLMs during implementation, thereby preventing confusion or ineffective execution, and it facilitates debugging of the entire pipeline by making the LLMs’ reasoning process transparent and traceable.
-Our ablation study in § 5.4.2 demonstrates that having a plan leads to better performance than not having one, consistent with findings from other domain applications [27].
-Figure 5: Prompt template for plan synthesis.
-To synthesize the implementation plan for the checker, we have designed an LLM-based agent whose prompt template is shown in Figure 5.
+A single bug pattern identified from a patch can be expressed with varying scope and complexity.
+Consider the Null-Pointer-Dereference involving devm_kzalloc (Figure 2(a)).
+A broad pattern (e.g., check any potentially null return) is comprehensive, but identifying all relevant functions/conditions poses significant static analysis challenges, hindering robust implementation by LLMs.
+Consequently, our approach favors more targeted bug patterns derived from the patch context.
+These facilitate precise and tractable checker synthesis by the LLMs.
+For the devm_kzalloc example, focusing specifically on its return value yields a targeted pattern that effectively addresses the observed bug class while being significantly more manageable for the LLM to implement correctly compared to the broader, more complex alternative.
+3.1.2. Plan Synthesis
+Once the bug pattern is identified, KNighter generates a high-level plan for implementing the static analyzer.
+This plan serves two critical purposes: first, it provides structured guidance to the LLMs during implementation, preventing confusion and promoting effective execution.
+Second, it facilitates debugging of the entire pipeline by making the LLMs’ reasoning process transparent and traceable.
+Our ablation study in § 5.4.2 confirms the value of this plan synthesis, demonstrating improved performance consistent with findings in other domains (sun2023pearl, ).
+For instance, synthesizing a checker for the unchecked devm_kzalloc return value pattern (illustrated in Figure 2(c)) might generate a plan with key steps such as: (1) Using program state to track memory regions from devm_kzalloc, (2) monitoring conditional branches (checkBranchCondition) to mark regions as checked if a null check occurs, and (3) detecting uses (checkLocation) of unchecked regions, potentially signaling a bug.
+This high-level structure guides the subsequent implementation phase.
+To synthesize the implementation plan for the checker, we have designed an LLM-based agent whose prompt template is shown in Figure 5(b).
 This agent takes the previously summarized bug pattern as input.
@@
 By including the signatures and brief descriptions of these utility functions in the prompt, we enable LLMs to leverage them effectively during the planning process, simplifying the overall task.
-3.1.3 Analyzer Implementation and Syntax Repair
-After identifying the bug pattern and making the plan, we leverage an LLM-based agent to implement the corresponding checker, as illustrated in Figure 6.
+(a) Prompt template for bug pattern analysis
+(b) Prompt template for plan synthesis.
+Figure 5. Prompt templates used by KNighter.
+3.1.3. Implementation and Syntax Repair
+After identifying the bug pattern and making the plan, we leverage an LLM-based agent to implement the corresponding checker.
 To maximize implementation accuracy, we provide the agent with comprehensive inputs: the distilled bug pattern and a structured implementation plan.
-We also provide a pre-defined checker template that standardizes the implementation structure and reduces potential errors.
+We also provide a pre-defined checker template, as shown in Figure 6, which standardizes the implementation structure and reduces potential errors.
 Moreover, we provide a list of utility functions that could help with the implementation.
-To handle potential compilation errors in the generated checkers, we employ a dedicated debugging agent.
-Inspired by existing work on program repair [32], this agent automatically processes compiler error messages and applies necessary fixes, effectively addressing syntax errors that may arise from LLM hallucinations.
+The synthesized checkers could have compilation errors, e.g., using the wrong static analysis API or incorrect variable types.
+To handle the potential compilation errors, we employ a dedicated debugging agent.
+Inspired by existing work on program repair (xia2024automated, ), this agent automatically processes compiler error messages and applies necessary fixes, effectively addressing syntax errors that may arise from LLM hallucinations.
 This automated debugging pipeline ensures that the final checkers are both syntactically correct and compilable.
-Figure 6: Prompt template for analyzer implementation.
-3.1.4 Validation
-To validate our checkers semantically and mitigate hallucinations by LLMs, we evaluate them against both the buggy (pre-patch) and patched versions of the Linux kernel.
-This dual-version analysis verifies that the checker can both detect original bugs and recognize their fixes.
-For efficiency, we restrict the evaluation to only the files modified by the patch and their dependencies, rather than scanning the entire kernel codebase.
-This focused approach enables rapid validation while maintaining effectiveness.
-We assess checker validity by comparing the number of reported issues between versions, acknowledging that while similar bugs might persist in the patched version, an effective checker should yield a meaningful difference in detection results.
-Detailed implementation specifics are presented in § 4.
-3.2 Checker Refinement
-After synthesizing a valid checker, we scan the entire Linux codebase.
-While effective at identifying potential issues, these checkers may be overly conservative, flagging correct code as problematic and generating false positives.
-To address this limitation, we develop an iterative refinement process powered by LLMs.
-The process evaluates each bug report’s validity and uses false positive cases to refine the checker.
-However, this refinement process faces two significant challenges:
-  • Bug reports often contain extensive context and implementation details, making them difficult to process efficiently.
-  • Debugging and refining checkers based on false positive reports requires complex analysis and careful modification.
-To address these challenges, we first distill bug reports to their essential elements.
-We extract only the "relevant lines" identified by the LLVM report and the trace path for each potential issue—preserving critical diagnostic information while eliminating extraneous context.
-We then deploy a specialized LLM-based triage agent to classify each report as either a true or false positive.
-Note, our classification criteria focus specifically on pattern alignment: we evaluate whether each reported issue exhibits the same bug pattern as the original input patch, rather than conducting a general correctness assessment.
-For reports classified as false positives, a dedicated refinement agent analyzes the specific case and modifies the checker accordingly, enhancing precision while preserving the checker’s ability to detect the original vulnerability pattern.
+Figure 6. Pre-defined checker template for CSA.
+3.1.4. Validation
+To semantically validate our checkers and mitigate potential inaccuracies from LLMs, we evaluate them against both the buggy (pre-patch) and patched versions of the relevant code (e.g., Linux kernel files).
+This differential analysis verifies that a checker correctly identifies the target bug in the original code and confirms its absence after the patch.
+For efficiency, we scope this validation to only the files modified by the patch and their dependencies, rather than the entire codebase.
+A checker is considered valid if it flags the bug in the pre-patch version and shows a corresponding reduction or elimination of that specific warning in the patched version.
+More details are in § 4.
+Figure 7. A report labeled as FP by our triage agent.
+3.2. Checker Refinement
+Following synthesis, each valid checker is used to scan the entire system.
+However, its initial validation doesn’t prevent potential false positives when analyzing the broader codebase, where correct code might be flagged erroneously.
+To mitigate this, we implement an iterative refinement procedure driven by LLMs.
+This involves evaluating the generated bug reports and feeding identified false positives back to refine the checker.
+However, automating this refinement faces hurdles.
+First, bug reports are often verbose, containing extensive context that is difficult to process efficiently.
+Second, debugging the checker logic and modifying it correctly based on false positives requires complex analysis.
+Our refinement pipeline addresses these challenges methodically.
+First, to manage report complexity, we distill generated bug reports to their essential components—primarily the “relevant lines” highlighted by the static analyzer (e.g., CSA (clang-static-analysis, )) and the corresponding trace path—stripping extraneous context while preserving critical diagnostics.
+Second, to navigate the complexity of analysis and modification, we employ specialized LLM-based agents.
+A triage agent classifies each distilled report, focusing strictly on pattern alignment (i.e., does the report match the original input bug’s pattern?) rather than general code correctness.
+If the triage agent identifies a report as a false positive, as exemplified in Figure 7, a dedicated refinement agent takes over.
+In the case shown, the initial checker (derived from the patch in Figure 2(a)) flagged the use of pmx->pfc because its logic failed to recognize if (unlikely(!pmx)) as a valid null check, perhaps confused by the unlikely() macro.
+The triage agent correctly interprets the check semantically and flags the report as FP.
+The refinement agent then uses this information to adjust the checker’s logic, specifically enhancing its ability to handle constructs like unlikely(), thereby preventing this type of false positive in subsequent scans while ensuring it can still detect the original vulnerability.
 A refined checker is accepted only if it satisfies two criteria: (1) it no longer generates warnings for the previously identified false positive cases, and (2) it maintains its validity by correctly differentiating between the original buggy and patched code versions.
 This criterion ensures the semantic accuracy of the refined checkers.
-4 Implementation
+4. Implementation
 Input commit collection.
-To collect patch commits, we implemented a systematic classification and selection process.
+To collect patch commits for rigorous evaluation, we implemented a systematic classification and selection process.
 First, we established 10 distinct bug categories.
@@
 Table 1 presents our categorization of 10 bug types and their corresponding patch commit counts.
-Table 1: Distribution of patch commits across 10 bug categories and the validity status of their synthesized checkers.
+These manually collected and labeled commits served as our benchmark dataset for rigorous evaluation.
+Table 1.
+Distribution of patch commits across 10 bug categories and the validity status of their synthesized checkers.
 “NPD” denotes “Null-Pointer-Dereference” and “UBI” indicates “Use-Before-Initialization”.
@@
 Few-shot examples.
-We prepared three end-to-end examples for LLMs’ in-context learning.
+We prepared three end-to-end examples for in-context learning.
 These three are patch commit 3027e7b15b02 (Null-Pointer-Dereference), 3948abaa4e2b (Use-Before-Initialization), and 4575962aeed6 (Double-Free).
 The design and implementation of the checker for these three commits required approximately 40 person-hours.
+This was a one-time effort, yielding reusable examples.
+We also explore the use of real-world, off-the-shelf examples (§ 5.4.2).
 Utility functions.
-While implementing checkers for the example commits, we identified several commonly needed utility functions.
-We implemented 9 utility functions, including getMemRegionFromExpr, to support checker development.
-These utility functions are designed to be simple to implement and extend.
+While implementing example checkers, we identified several common helper operations.
+We implemented 9 such utility functions (e.g., getMemRegionFromExpr) to encapsulate low-level Clang Static Analyzer tasks, simplifying checker development, particularly for LLM synthesis.
+These utilities were designed for simplicity and extensibility.
 Valid checkers.
@@
 Next, we check out the repository to the buggy commit (immediately preceding the patch) and scan these objects to count the number of bug reports ($N_{buggy}$).
-We then scan the same objects after applying the patch commit to obtain the number of remaining bug reports ($N_{patched}$).
-A checker is considered valid if $N_{buggy}>N_{patched}$ and $N_{patched}<T_{valid}$, where $T_{valid}$ is a threshold value, which is 50 by default.
+We then scan these objects after applying the patch commit to obtain the number of remaining bug reports ($N_{patched}$).
+A checker is considered valid if $N_{buggy}>N_{patched}$ and $N_{patched}<T_{valid}$, where $T_{valid}$ is a threshold value (50 by default).
 Plausible checkers.
@@
 For checkers failing these criteria, we implement an iterative refinement protocol targeting the identified false positives, permitting up to three refinement iterations to improve precision.
-5 Evaluation
+5. Evaluation
 We explore the following research questions for KNighter:
   1. RQ-1.
-    Can KNighter synthesize high-quality static analyzers?
+    Can KNighter generate high-quality checkers?
   2. RQ-2.
@@
   3. RQ-3.
-    How do the static analyzers synthesized by KNighter compare with human-written ones?
+    Are the capabilities of KNighter orthogonal to the human-written checkers?
   4. RQ-4.
@@
   • Compilation Failures: Checkers that fail during compilation due to syntax or dependency errors.
-  • Runtime Errors: Checkers that compile successfully but crash during execution (e.g., "The analyzer encountered problems on source files").
+  • Runtime Errors: Checkers that compile successfully but crash during execution (e.g., ”The analyzer encountered problems on source files”).
   • Semantic Issues: Checkers that cannot distinguish between the buggy and patched code.
+Static Analysis Capabilities. We further examine the static analysis capabilities employed by checkers, including path sensitivity, region sensitivity, and advanced state tracking.
 Hardware and software.
@@
 By default, when scanning the entire codebase, we use -j32.
-We evaluated using Linux v6.13, and for bug finding, we examined versions from v6.9 to v6.13.
+We evaluated using Linux v6.13, and for bug finding, we examined versions from v6.9 to v6.15.
 The Linux configuration used is allyesconfig.
-5.1 RQ1: Synthesized Checkers
-(a) The bug in the Use-Before-Initialization patch.
-(b) A false positive reported by KNighter.
-Figure 7: Reported false positive for a UBI commit.
-Synthesis.
-From the commits collected in Table 1, we successfully generated valid checkers for 39 commits.
-Furthermore, KNighter successfully generated valid checkers across diverse bug types beyond those in our few-shot examples, demonstrating the generalizability of our approach.
-The complete synthesis process required 15.9 hours, consuming 8.2 million input tokens and producing 1.2 million output tokens.
-For commits yielding valid checkers, KNighter required an average of 2.4 attempts per checker (maximum: 8 attempts).
-The generated valid checkers averaged 125.7 lines of code.
-Refinement.
+5.1. RQ1: Synthesized Checkers
+(a) Bug in Use-Before-Initialization patch.
+(b) False positive report for UBI.
+Figure 8. Examples of false positives by KNighter.
+We evaluate KNighter on the 61 commits listed in Table 1 to show that it can synthesize high-quality checkers across various bug types, beyond those in our few-shot examples.
+5.1.1. Checker Synthesis
+In total, valid checkers were generated for 39 commits.
+The 39 synthesized checkers average 125.7 lines of code and exhibit diverse static analysis capabilities.
+Specifically, 37 checkers are path-sensitive, 13 incorporate region sensitivity, and 16 employ advanced state tracking; in contrast, only 2 leverage AST travelers.
+This suggests that KNighter can generate complex analysis logic, not just simple pattern matching.
+Cost.
+The full synthesis process required 15.9 hours, during which 8.2 million input tokens were processed and 1.2 million output tokens produced, resulting in an approximate cost of \$0.24 per commit using O3-mini.
+For commits that ultimately yielded valid checkers, an average of 2.4 synthesis attempts was necessary (with a maximum of 8 attempts observed).
+Failure analysis. We now break down the failures from two perspectives: the underlying failure root causes and the observed failure symptoms during the synthesis process.
+(i) Failure root causes. Among the 61 commits processed, 22 did not result in any valid checker. Our investigation into these failures indicates that:
+  • 2 (9%) commits failed due to an inaccurate bug pattern,
+  • 7 (32%) failed owing to an inaccurate plan, and
+  • 13 (59%) were caused by inaccurate implementation.
+For the implementation-related failures, a common issue was that compiler optimizations inlined certain function calls (e.g., strcp and memset), which prevented the checker from properly intercepting these calls.
+Our approach exhibits limitations in handling buffer overflow and use-after-free commits.
+We believe these challenges stem from two main factors: static analysis inherently struggles with precise value determination—especially when establishing buffer bounds at compile time—and it also faces significant hurdles in analyzing multi-threaded code, for instance, when assessing the proper use of locks.
+(ii) Failure symptoms. During synthesis (allowing up to 10 attempts per commit), 273 failed attempts were recorded from all 61 commits. The failures can be categorized as:
+  • 65 attempts (23.8%) resulted in compilation errors,
+  • 1 attempt (0.4%) led to a runtime error, and
+  • 207 attempts (75.8%) suffered from semantic issues that obstructed proper bug identification.
+Of the 207 semantic failures, 34 checkers erroneously flagged both buggy and patched code as potentially problematic, while the remaining 173 misclassified both versions as bug-free.
+This outcome underscores the challenge of accurately distinguishing buggy code.
+Interestingly, even the 173 checkers that did not recognize the specific bug in their input patch can still be valuable.
+When deployed across the large system, these checkers may successfully detect bugs with similar patterns in other contexts.
+This apparent paradox likely arises because the failure to detect the training bug is sometimes due to edge cases or context-specific complexities rather than inherent deficiencies in the checkers’ detection logic.
+Moreover, these checkers generally exhibit lower false positive rates compared to those that incorrectly flag both buggy and patched code, enhancing their practical utility for bug detection.
+5.1.2. Checker Refinement.
 After scanning the entire kernel codebase with these 39 valid checkers, 26 of them were labeled “plausible” directly.
@@
 In total, 19 refinement steps were completed successfully.
-This demonstrates the effectiveness of our refinement pipeline, which successfully refined 84.6% of the valid checkers that were not initially labeled as “plausible”.
+This demonstrates the effectiveness of our refinement pipeline, which successfully refined 84.6% of the valid checkers that were not “plausible” initially.
 False positive rate.
@@
 This indicates that the combination of our plausible checkers and bug triage agent has a false positive rate of 32.2%.
-Here are some example checkers with poor performance (high false positive rate).
-For the commit 90ca6956d383 (“ice: Fix freeing uninitialized pointers”) shown in Figure 7(a), the issue occurs when the pointer pcaps is not set to NULL, and an early return or error path is taken before it is assigned a valid allocation, causing the cleanup logic to attempt to free an uninitialized (or garbage) pointer.
-However, to prove such a bug pattern exists, there must be an early exit-path where the pointer is not assigned any value.
-Thus, for the bug report shown in Figure 7(b), although the cert pointer is not initialized, it will be directly assigned a value and there is no path where it remains unassigned.
-Therefore, this is not a bug.
-Our synthesized checker failed to consider this constraint, and our triage agent also overlooked this issue.
-Invalid checkers.
-During the checker synthesis process (up to 10 attempts for each commit), we recorded in total 273 failed attempts across all 61 commits.
-These invalid checkers failed for three primary reasons: 65 encountered compilation errors, 1 exhibited runtime errors, and the remaining 207 contained semantic issues that prevented proper bug identification.
-Of the 207 checkers with semantical issues, 34 incorrectly labeled both buggy and patched code as potential bugs, while the remaining 173 incorrectly labeled both versions as not containing bugs.
-This demonstrates that correctly identifying buggy code remains a significant challenge for KNighter.
-Interestingly, the 173 checkers that fail to identify the specific bug in their input patch can still provide value.
-When applied to the broader kernel codebase, these checkers may successfully detect bugs with similar patterns in other contexts.
-This apparent paradox occurs because a checker’s failure to detect its training bug might be due to specific edge cases or context-dependent factors rather than fundamental flaws in its detection logic.
-For instance, a checker might not detect a null pointer dereference in its training example due to complex control flow, but still correctly identify simpler instances of the same pattern elsewhere.
-Moreover, these checkers often have lower false positive rates compared to those that incorrectly flag both buggy and patched code as problematic, making them even more practical for real-world bug detection.
-Our approach shows limitations in detecting buffer overflow and use-after-free vulnerabilities.
-A key challenge is that CSA has inherent difficulty in performing precise value analysis, particularly in determining buffer bounds at compile time.
-This limitation makes it challenging to statically reason about potential overflow conditions that depend on runtime values.
-Table 2: Detected bugs by KNighter.
+Our manual analysis of the 29 false positives revealed three recurring patterns leading to incorrect reports:
+  • Inaccurate bug pattern: In 5 cases, although the checker correctly identified the original bug/patch scenario, the inferred bug pattern lacked the necessary precision for reliable detection across different contexts.
+  • Incorrect pattern matching: For 6 reports, the checker correctly identified the bug pattern but applied it too broadly, flagging code segments that did not meet the specific constraints intended by the pattern.
+  • Trigger condition mismanagement: The most common issue (18 reports) involved checkers where both the pattern and matching were correct, but the checker failed to manage trigger or state conditions properly (e.g., failing to recognize a pointer had already been validated before use).
+Case study: A high false positive rate checker.
+In commit 90ca6956d383 (“ice: Fix freeing uninitialized pointers”, see Figure 8(a)), the issue stems from the pointer pcaps not being initialized to NULL.
+If an early return or error path occurs before the pointer is allocated, the cleanup routine may inadvertently attempt to free an uninitialized (or garbage) pointer.
+In such cases, the checker should account for the possibility of an early exit leaving the pointer unset.
+In contrast, the bug report in Figure 8(b) highlights a scenario where, although the cert pointer starts uninitialized, it is immediately assigned a valid value along every execution path, ensuring it is never left in an unassigned state.
+Thus, despite the initial uninitialized state, the code does not constitute a bug.
+Our synthesized checker did not incorporate these nuanced constraints, and the triage agent likewise failed to recognize the critical differences, ultimately leading to a false positive.
 Total Confirmed Fixed Pending CVE
-KNighter 70 56 41 15 11
-(a) Number of bugs in each subsystem.
+KNighter 92 77 57 15 16
+(a) Bug summary statistics detected by KNighter.
 (b) Number of bugs in each type.
-Figure 8: The details of newly detected bugs by KNighter.
+(c) Number of bugs in each subsystem.
+(d) Number of bugs with different lifetimes.
+Figure 9. The details of newly detected bugs by KNighter. Subfigure (9(a)) summarizes overall status, while (9(b)), (9(c)), and (9(d)) show breakdowns by type, subsystem, and lifetime.
+5.2. RQ2: Detected Bugs
+5.2.1. Overall.
+To date, static analyzers synthesized by KNighter have identified 92 new bugs in the Linux kernel.
+As summarized in Figure 9(a), developer confirmation has been received for 77 of these bugs.
+Among those confirmed, 57 have already been fixed.
+The remaining 15 bugs are currently awaiting developer review (calculated as Total - Confirmed).
+Notably, 16 of the discovered bugs have been assigned CVE numbers, showing their practical security impact.
 (a) Input Use-After-Free patch.
@@
 (d) CVE-2024-50259, found by the checker for the patch above.
-Figure 9: Example vulnerabilities detected by KNighter.
-5.2 RQ2: Detected Bugs
-To date, the static analyzers generated by KNighter have detected 70 new bugs, with 56 confirmed and 41 already fixed (see Table 2).
-Notably, 11 of these bugs have been assigned CVE numbers.
-We identified more bugs than those reported as true positives in RQ-1 because our analysis spanned a broader range of Linux kernel versions.
-Figure 8 shows the details of newly detected bugs, including their subsystem locations and bug types.
-We found that most bugs are from the drivers subsystem (45/70), which aligns with the fact that driver code constitutes the largest portion of the kernel [6].
-Following drivers, we also found 12 and 6 bugs in the sound and net subsystems respectively.
-As shown in Figure 8(b), KNighter demonstrates capability in detecting a diverse range of bug types.
-Null-Pointer-Dereference issues constituted the largest category (30/70 bugs), followed by Integer-Overflow and Misuse vulnerabilities.
-This distribution likely reflects the pattern-based nature of Null-Pointer-Dereference and Misuse bugs, which makes their checkers more amenable to LLM-based implementation.
-Integer-Overflow bugs, typically involving variable type mismatches, benefit from the Clang toolchain’s strong type analysis capabilities.
-More complex issues such as Use-After-Free and Concurrency bugs proved more challenging to detect at compile time, particularly those involving multi-threaded execution contexts.
-Overall, KNighter-synthesized checkers are general to detect various types of bugs.
-Here are example vulnerabilities detected by KNighter.
+Figure 10. Example vulnerabilities detected by KNighter.
+The checkers for bug detection originate from two sources: (i) the initial 61 manually collected commits used for evaluation in § 3.1 across diverse bug types (shown in Table 1), and (ii) an additional set of 100 commits automatically collected using keywords related to Null-Pointer-Dereference to further explore that specific bug class.
+Figure 9(b) and Figure 9(c) show the bug distribution from these sources—with the light blue for bugs from the initial evaluation set and the light purple for those from the auto-collected set.
+Bug types.
+Analysis of checkers from the manually collected commits reveals that KNighter can find a wide range of bug types (in Figure 9(b)).
+Among these, Null-Pointer-Dereference bugs are the most prevalent, showing the KNighter’s strength in this area.
+In response, we expanded our effort by automatically collecting commits related to Null-Pointer-Dereference, which yielded an additional 30 bugs.
+Bug location.
+As shown in Figure 9(c), the detected bugs span various Linux kernel subsystems.
+The majority appear in the drivers subsystem (62 out of 92), reflecting its large footprint in the kernel (bursey2024syzretrospector, ).
+Additionally, 14 and 7 bugs were identified in the sound and net subsystems, respectively.
+Notably, 2 bugs were found in the samples directory—an area that provides example usage for kernel developers and where correctness is especially critical.
+Bug lifetime.
+Figure 9(d) illustrates the distribution of bug lifetimes.
+Notably, the average bug lifetime is 4.3 years, and 26 bugs had existed for over five years before detection.
+This indicates that the vulnerabilities uncovered by KNighter are difficult to detect and remain latent for extended periods, underscoring the effectiveness of our approach.
+5.2.2. Case Study.
+Here are examples of vulnerabilities detected by KNighter.
 CVE-2025-21715.
@@
 CVE-2024-50259.
-Figure 9(c) shows an input patch addressing a buffer overflow vulnerability.
+Figure 10(c) shows an input patch fixing a buffer overflow vulnerability.
 The patch mitigates the risk by limiting the number of bytes copied via copy_from_user to sizeof(mybuf) - 1, thereby preserving space for a trailing zero.
 This trailing zero is essential for subsequent string operations, such as sscanf, to function correctly.
-Leveraging this patch as a reference, the checker generated by KNighter identified a similar bug in nsim_nexthop_bucket_activity_write, as shown in Figure 9(d).
+Taking this patch as input, the checker generated by KNighter identified a similar bug in nsim_nexthop_bucket_activity_write, as shown in Figure 10(d).
 In this case, the omission of appending a trailing zero after copying data from userspace could lead to improper string handling and potential overflow issues.
 This detected issue was subsequently fixed by adding the trailing zero and was assigned CVE-2024-50259.
-5.3 RQ3: Baseline Comparison
+5.3. RQ3: Orthogonality with Smatch
 Since no comparable automated static analyzer generation approaches exist for this domain, we evaluate KNighter against expert-written checkers.
-Our baseline is provided by Smatch [1], which is widely used in Linux kernel analysis and supports tailored checks for all bug types considered in Table 1.
+Our baseline is provided by Smatch (smatch, ), which is widely used in Linux kernel analysis and supports tailored checks for all bug types considered in Table 1.
 We conducted the comparative analyses by running Smatch on the entire codebase to determine if it could detect the bugs found by KNighter.
-Our evaluation revealed that Smatch reported 1970 errors and 2870 warnings across the kernel.
-Notably, it failed to detect any of our true positive bugs, underscoring KNighter’s unique detection capabilities.
-Further investigation of Smatch’s checkers showed they don’t fully leverage domain-specific knowledge available in the Linux kernel—knowledge that KNighter extracts from historical patches.
+Smatch reported a total of 1970 errors and 2870 warnings across the kernel.
+We manually inspected all the files where bugs were detected to assess whether any of the true positive bugs identified by KNighter were also detected by Smatch.
+Notably, Smatch failed to detect any of our true positive bugs, underscoring the unique detection capabilities of KNighter.
+Further analysis of Smatch’s checkers revealed that they do not fully leverage the domain-specific knowledge embedded in the Linux kernels—a resource that KNighter effectively extracts from historical patches.
 For instance, Smatch’s check_deref checker employs static range analysis to identify potential null pointers but lacks domain-specific insights.
@@
 We conclude that KNighter and Smatch detect different classes of bugs, demonstrating KNighter’s effectiveness in learning domain knowledge from patches and subsequently identifying diverse bugs and vulnerabilities.
-5.4 RQ4: Effectiveness of Components
-5.4.1 Bug Triage Agent
-To evaluate our triage agent, from the valid checkers, we sampled up to 5 reports per checker to reduce manual inspection efforts.
-In total, we collected 79 reports from 18 checkers.
-The remaining 21 valid checkers did not identify any bugs, while 2 non-plausible checkers also generated reports.
-These reports may include those generated during the refinement process.
-Our objective here was to evaluate the effectiveness of our triage agent.
-Among these reports, 29 were labeled as “bug” (positive) by our triage agent, while the remaining 50 were labeled as “not-a-bug” (negative).
-Two of our authors manually cross-checked all reports to establish ground truth.
-Our triage agent achieved 7 true positives (TP), 22 false positives (FP), 50 true negatives (TN), and zero false negatives (FN).
-Notably, the absence of false negatives is beneficial because it helps ensure that most true bugs are captured, even if this comes with some false positives that can be filtered out in subsequent steps or human inspection.
-5.4.2 Ablation Study
-Table 3: Ablation study results.
-“Default” means KNighter’s standard configuration utilizing multi-stage synthesis and the O3-mini language model.
+5.4. RQ4: Effectiveness of Components
+5.4.1. Bug Triage Agent
+To evaluate our bug triage agent, from the 39 valid checkers, we sampled up to 5 reports per checker to reduce manual inspection efforts, aiming to reduce manual inspection efforts while maintaining evaluation coverage.
+In total, we collected 79 reports from 18 checkers, while the remaining 21 valid checkers didn’t generate reports.
+Our triage agent classified 29 reports as “bug” (positive) and 50 as “not-a-bug” (negative).
+These classifications were compared against ground truth labels established by manual review from two authors.
+The agent achieved 7 true positives (TP), 22 false positives (FP), 50 true negatives (TN), and zero false negatives (FN).
+The absence of false negatives is particularly important, as it indicates the agent effectively prioritized all potentially true bugs in this set, minimizing the risk of overlooking genuine issues, even though the 22 false positives require further filtering.
+5.4.2. Ablation Study
+Table 2.
+Ablation study results.
+“Default” means KNighter’s standard configuration utilizing multi-stage synthesis, fixed few-shot examples, and the O3-mini model.
 Alternative configurations are compared against this baseline.
@@
 W/o multi-stage 8 52 3 75
+W/ RAG 12 37 4 62
 W/ GPT-4o 11 31 0 76
+W/ DeepSeek-R1 11 29 8 66
 W/ Gemini-2-flash 4 130 2 44
-To further evaluate our design choices, we created a sample dataset of patch commits for an ablation study.
+To evaluate our design choices, we created a sample dataset of patch commits for an ablation study.
 We randomly sampled 2 commits from each bug type using zero as the random seed.
 This resulted in a dataset of 20 commits (2 commits × 10 bug types).
-Table 3 shows the results of our ablation study.
+Table 2 shows the overall results of our ablation study.
 Checker synthesis.
-Our default pipeline implements a three-stage approach: bug pattern analysis, detection plan synthesis, and checker implementation.
-We evaluated the effectiveness of directly synthesizing checkers without the initial bug pattern analysis and plan synthesis phases while maintaining identical in-context learning examples across configurations.
-As shown in Table 3, without multi-stage synthesis, KNighter produced valid checkers for only 8 commits compared to 12 with our default approach.
-The single-stage approach also generated significantly more syntax errors (52 vs.
-28), resulting in checkers that failed to compile.
+First, we assessed the impact of our default three-stage synthesis approach (bug pattern analysis, plan synthesis, checker implementation) compared to directly synthesizing checkers in a single stage (omitting explicit pattern/plan steps), using identical few-shot examples.
+As shown in Table 2, the multi-stage approach proved more effective, yielding valid checkers for 12 commits compared to only 8 for the single-stage method.
+Furthermore, the single-stage approach resulted in significantly more syntax errors (52 vs.
+28), often leading to checkers that failed to compile.
+This highlights the value of the structured multi-stage process for improving both validity and compilability.
+Second, we explored using Retrieval-Augmented Generation (RAG) (gao2023retrieval, ) for selecting few-shot examples dynamically, comparing it against our default set of fixed examples.
+It utilizes a knowledge base derived from 118 official CSA checkers (clang-static-analysis, ), embedded using text-embedding-ada-002 (text-embedding, ).
+During synthesis, three relevant examples are retrieved based on semantic similarity.
+Our results indicate that RAG-based example selection achieved comparable effectiveness to our fixed examples, also generating valid checkers for 12 commits.
+However, because the official CSA checkers used are substantially longer than our curated fixed examples, the RAG approach incurred approximately double the input/output token cost.
+Due to similar effectiveness, our default fixed few-shot examples offer better cost-efficiency for this task.
 LLM choice.
-We evaluated KNighter’s performance across different language models for checker synthesis.
-In addition to our default model O3-mini (reasoning model), we tested GPT-4o and Gemini-2-flash.
-As Table 3 demonstrates, GPT-4o performed slightly worse than O3-mini, generating 11 valid checkers compared to 12.
-Gemini-2-flash performed substantially worse, producing valid checkers for only 4 commits.
-Upon closer inspection, we found that Gemini-2-flash struggled with Clang Static Analyzer implementation, frequently using non-existent APIs and generating syntax errors at a much higher rate (130 vs.
+We evaluated the performance of KNighter across different language models for checker synthesis.
+In addition to our default model, O3-mini, we tested GPT-4o, Gemini-2-flash, and open-source DeepSeek-R1.
+As detailed in Table 2, O3-mini yielded the most valid checkers (12).
+GPT-4o and DeepSeek-R1 performed comparably to each other, generating 11 valid checkers each, only slightly fewer than O3-mini.
+This suggests that multiple high-capability models, including open-source options, are viable, although minor differences in performance exist.
+In contrast, Gemini-2-flash performed substantially worse, producing valid checkers for only 4 commits.
+Upon closer inspection, we found that Gemini-2-flash struggled with CSA implementation, frequently using non-existent APIs and generating syntax errors at a much higher rate (130 vs.
 28).
-6 Related Work
-6.1 Traditional Static Analysis
-Given the cruciality of Linux kernel and the diversity of its vulnerabilities, many static analyzers have been developed to target different classes of bugs.
-CRIX [20] focuses on detecting missing-check bugs in kernel by constructing def-use slices of critical variables.
-Goshawk [21] proposes a structure-aware and object-centric abstraction over memory management operations to aid detecting double-free and use-after-free bugs.
-UBITect [39] integrates type qualifier analysis and symbolic execution to detect use-before-initialization bugs in the Linux kernel, which might encounter challenges due to exhaustive path exploration.
-CRed [34] detects Use-After-Free bugs in scale by spatio-temporal context reduction.
-LR-Miner [16] establishes field-sensitive locking rules between locks and variables.
-Then it detects rule violations to pinpoint data races in mainstream OS kernels.
-Authors of [12] extensively analyze over one thousand reference counting bugs in the history of Linux kernel and derived 9 anti-patterns.
-They further developed a set of static analyzers to detect such patterns.
-LinkRID [19] also targets refcounting bugs by modelling common Linux-specific refcount usage conventions.
-DCUAF[5] conducts local-global analysis to find concurrent functions to discover kernel currency bugs.
-SUTURE[41] performs high-order taint analysis to discover kernel vulnerabilities caused by malicious userspace input.
-While traditionally designed static analyzers can perform well on their targeted classes of bugs, they require extensive analysis and implementation effort from human experts.
-This limits their ability to scale to additional bug classes and keep pace with the rapid evolution of the Linux kernel.
-KNighter is complementary to such analyzers by detecting more classes of bugs and adapting to newly developed bug patterns.
-Beyond one specific class of bugs, many static analysis systems implement different analysis techniques to improve precision and efficiency.
-FiTx [28] implements fast analysis for single-compilation unit without the requirement of complex features like indirect call analysis.
-PATA [17] improves the alias analysis precision using control-flow paths and access paths information.
-FiTx and PATA support detecting 6 and 3 common bug patterns respectively, by encoding bug patterns as finite state machines.
-While these static analysis systems can potentially support more classes of bugs, analyzing bugs and designing the pattern state machines remain challenging and labor-intensive.
-KNighter automates this process by automatically synthesizing bug patterns from patches.
-Beyond automation, although KNighter generates CSA checkers, the idea is generalizable to using other static analysis tools without fundamental obstacles since LLMs have been trained on most existing tools.
-6.2 LLM-Based Static Analysis
-With the recent advances of LLMs, many techniques leverage LLMs’ code comprehension ability to mediate the limitations of traditional static analysis tools that typically require approximation or expert-produced knowledge.
-IRIS [18] enhanced CodeQL’s taint analysis by using LLMs to infer taint specifications for external APIs, which would have required manual labeling.
-To detect Use-Before-Initialization bugs, LLift [15] uses LLMs to extract initializers and generate their post-constraints.
-The post-constraints are then used during further analysis to prune unreachable paths.
-InferROI [30] uses LLMs to infer resource-oriented intentions and then perform lightweight static resource leak detection.
-While the above approaches can enhance static analyzers, a substantial portion of the analyzer still needs to be crafted by researchers.
-In contrast, KNighter can automatically synthesize the entire checker without human intervention.
-Beyond this, the checkers generated by KNighter could potentially be further complemented by additional information provided by the above approaches.
-Another branch of work directly use LLMs to analyze the code without traditional static analysis tools.
-Vul-RAG [10] constructs a knowledge base from existing CVE instances and directly queries LLMs with Retrieval-Augmented-Generation to determine whether a new piece of code has the same bug.
-Zhang et al.
-[40] studies various prompt design techniques for vulnerability detection.
-Fine-tuning has also been explored by [25] to perform vulnerability detection.
-The above LLM-Based analysis techniques require querying LLMs with the code under test.
-In the context of Linux kernel, querying LLMs with millions lines of code, potentially multiplied by the number of different classes of bugs, is prohibitively expensive.
-In comparison, KNighter generates a static analysis checker with negligible token cost that can be used to efficiently scan the kernel codebase.
-7 Conclusion
+This highlights a crucial insight: successful checker synthesis demands more than general coding proficiency; accurate knowledge or inference of the target framework’s specific APIs and conventions (CSA in this case) is essential.
+6. Related Work
+6.1. Traditional Static Analysis
+Given the cruciality of the Linux kernel and the diversity of its bugs, many static analyzers have been developed. These generally fall into several categories based on their approach.
+Rule/Model-based analyzers.
+A significant body of work focuses on detecting specific classes of bugs using predefined rules or models.
+Examples include CRIX (kernel-static-lu2019detecting, ) (detecting missing checks via def-use slices), Goshawk (kernel-static-lyu2022goshawk, ) (memory corruption analysis), UBITect (zhai2020ubitect, ) (Use Before Initialization bugs), CRed (yan18uaf, ) (Use-After-Free detection), LR-Miner (li2024lr, ) (data races via locking rules), tools targeting refcounting bugs based on derived anti-patterns (kernel-static-refcounting, ) or specific conventions (liu22linkrid, ), DCUAF (bai19uaf, ) (concurrent Use-After-Free), and SUTURE (zhang21suture, ) (taint analysis for userspace input vulnerabilities).
+While often effective for their targeted bug classes, these approaches typically require extensive human expertise for analysis design and implementation, limiting their scalability and adaptability to newly emerging bug patterns in the rapidly evolving kernel.
+Patch-based specification inference.
+Another line of work leverages historical patches to infer specifications, which are then used by separate checkers.
+For instance, APHP (aphp, ) extracts API Post-Handling (APH) specifications from both code and descriptions in patches to detect APH violations.
+A very recent work, Seal (seal, ), analyzes security patches to infer diverse specifications for Linux interfaces.
+While powerful, these methods generally rely on existing static analysis infrastructure to enforce the inferred specifications.
+We attempted to include Seal in our experimental comparison, but encountered practical difficulties as its publicly available version depends on private commercial tools and was not compatible with the recent Linux kernel versions.
+General static analysis frameworks.
+Other systems focus on improving underlying static analysis techniques for better precision or efficiency across multiple bug types.
+FiTx (suzuki2024fitx, ), for example, implements fast analysis for single compilation units, while PATA (li22pata, ) enhances alias analysis precision using path information.
+These systems often encode bug patterns as state machines (suzuki2024fitx, ; li22pata, ), but designing these patterns still typically involves manual effort.
+Our synthesis approach.
+In contrast to the above, our work, KNighter, uses LLMs to synthesize the static analyzer itself directly from historical patch information.
+Instead of relying on predefined rules or inferring specifications for existing checkers, KNighter learns both the bug pattern and the corresponding detection logic, generating a new checker automatically.
+This approach fundamentally differs by automating the analyzer creation process, potentially offering greater adaptability to new bug types compared to traditional methods or specification inference.
+This distinction may also make KNighter more versatile; while specification-inference tools excel at finding bugs related to the interfaces they model, KNighter can synthesize checkers for a broader range of issues demonstrated in patches, including problems not strictly tied to interface misuse, such as certain integer overflows or complex logic errors.
+Furthermore, although KNighter currently generates CSA checkers, the core idea of LLM-based checker synthesis is potentially generalizable to target other static analysis frameworks, given the broad training data of modern LLMs.
+Thus, KNighter complements existing approaches by providing a mechanism to automatically generate analyzers for newly observed bug patterns found in patches.
+6.2. LLM-Based Static Analysis
+Recent LLM advancements enable new static analysis techniques, differing significantly from our synthesis approach.
+LLM-augmented static analysis.
+Some techniques use LLMs to assist existing tools by automating tasks previously requiring manual effort.
+For example, LLMs can infer taint specifications for external APIs (IRIS (li2024llm, )), generate post-constraints to prune analysis paths (LLift (li2024enhancing, )), or infer resource-handling intentions (InferROI (wang2024resourceleak, )).
+While helpful, these methods still fundamentally depend on a substantial, human-developed analyzer core.
+This can limit their ability to easily adapt or generalize to detect novel bug types beyond those originally targeted by the core analyzer’s design.
+Direct code analysis with LLMs.
+Other work uses LLMs to directly analyze source code, using techniques like Retrieval-Augmented Generation (Vul-RAG (du2024vul, )), prompt engineering (cy24prompt, ), or fine-tuning (shestov, ).
+However, directly applying LLMs to scan large systems like the Linux kernel is often prohibitively expensive and faces scalability challenges.
+Our synthesis approach.
+In contrast, KNighter uses LLMs to synthesize the entire static analyzer from patches.
+Unlike LLM-augmented tools, this minimizes reliance on pre-existing manual analyzer development, enhancing adaptability for diverse bug types.
+Unlike direct LLM analysis, KNighter generates efficient, reusable static checkers, avoiding the high costs and scalability issues of scanning massive systems directly with LLMs.
+This makes KNighter a practical approach for applying LLM intelligence to large-scale static analysis, complementing other methods by automatically creating analyzers for new bug patterns.
+Checkers generated by KNighter could also potentially benefit from integration with LLM-augmented techniques.
+7. Conclusion
 This paper introduces KNighter, a novel approach that transforms how LLMs can contribute to static analysis for complex systems like the Linux kernel.
-By synthesizing specialized static analyzers rather than directly analyzing code, KNighter bridges the gap between LLMs’ reasoning capabilities and the practical constraints of analyzing massive codebases.
-KNighter’s practical impact is demonstrated by the discovery of 70 previously unknown bugs in the Linux kernel, with 56 confirmed and 41 already fixed.
-Looking forward, KNighter opens new possibilities for practical LLM-based static analysis.
+By synthesizing static analyzers rather than directly analyzing code, KNighter bridges the gap between LLMs’ reasoning capabilities and the practical constraints of analyzing massive systems.
+KNighter’s practical impact is shown by the discovery of 92 new, long-latent bugs in the Linux kernel, with 77 confirmed, 57 fixed, and 16 CVE assigned.
+Looking forward, KNighter opens new possibilities for scalable LLM-based static analysis.
 Future work could extend this approach to other systems beyond the Linux kernel, incorporate additional learning paradigms, and further refine checker generation techniques to address more complex bug patterns.

### v3

@@
 Our evaluation on the Linux kernel demonstrates that KNighter generates high-precision checkers capable of detecting diverse bug patterns overlooked by existing human-written analyzers.
-To date, KNighter-synthesized checkers have discovered 92 new, critical, long-latent bugs (average 4.3 years) in the Linux kernel; 77 are confirmed, 57 fixed, and 16 have been assigned CVE numbers.
+To date, KNighter-synthesized checkers have discovered 92 new, critical, long-latent bugs (average 4.3 years) in the Linux kernel; 77 are confirmed, 57 fixed, and 30 have been assigned CVE numbers.
 This work establishes an entirely new paradigm for scalable, reliable, and traceable LLM-based static analysis for real-world systems via checker synthesis.
+Static Analysis, Large Language Models
+journalyear: 2025copyright: ccconference: ACM SIGOPS 31st Symposium on Operating Systems Principles; October 13–16, 2025; Seoul, Republic of Koreabooktitle: ACM SIGOPS 31st Symposium on Operating Systems Principles (SOSP ’25), October 13–16, 2025, Seoul, Republic of Koreadoi: 10.1145/3731569.3764827isbn: 979-8-4007-1870-0/2025/10ccs: Security and privacy Systems securityccs: Software and its engineering Automated static analysis
 1. Introduction
@@
 We implement our approach in a tool, KNighter, the first fully automated pipeline for synthesizing static analyzers, built upon the open-source Clang Static Analyzer (CSA) (clang-static-analysis, ).
-Although the methodology generalizes to different systems, our evaluation targets the Linux kernel, one of the most fundamental software systems.
-From 61 historical patch commits, KNighter successfully synthesizes 37 high-quality checkers with a false positive rate of about 35%.
+While the methodology generalizes to different systems, we target the Linux kernel, one of the most fundamental software systems.
+In the evaluation of 61 diverse bug-fix patches, KNighter synthesized the high-quality checkers for 61% of them, achieving a false positive rate of about 35% aided by the triage agent.
+Demonstrating practical impact, KNighter has uncovered 92 new, long-latent vulnerabilities (average 4.3 years) in the Linux kernel, resulting in 77 developer confirmations, 57 fixes, and 30 CVEs.
 Furthermore, the vulnerabilities detected are orthogonal to those found by existing expert-written analyzers (smatch, ).
-Demonstrating practical impact, our synthesized checkers have uncovered 92 new, long-latent vulnerabilities (average 4.3 years) in the Linux kernel, resulting in 77 developer confirmations, 57 fixes, and 16 CVEs.
 These findings validate our approach’s efficacy and its contribution to system reliability.
@@
   • Approach. We implement KNighter with multi-stage synthesis and automated refinement pipelines for the Linux kernel. This design enables detection of diverse bug classes in large-scale systems.
-  • Evaluation.
-We evaluate KNighter by synthesizing checkers for the patch commits spanning multiple bug categories.
-Our results demonstrate that KNighter generates high-precision checkers with practical false positive rates.
+  • Evaluation. We demonstrate that KNighter successfully synthesizes effective checkers from the Linux kernel bug-fix patches across various bug categories, achieving practical false positive rates.
   • Real-world impact.
-KNighter-generated checkers have discovered 92 new, long-latent (average 4.3 years) bugs in the Linux kernel, with 77 confirmed, 57 fixed, and 16 assigned CVE numbers—demonstrating its practical impact on system reliability and security.
+KNighter-generated checkers have discovered 92 new, long-latent (average 4.3 years) bugs in the Linux kernel, with 77 confirmed, 57 fixed, and 30 assigned CVE numbers—demonstrating its practical impact on system reliability and security.
 KNighter is available at ise-uiuc/KNighter.
@@
 It updates another custom state map, PtrAliasMap, to track potential aliases between memory regions involved in the assignment, ensuring the checker correctly handles cases where multiple pointers might refer to the same potentially null memory.
+Figure 3. Overview of KNighter.
 2.2. Motivating Example
@@
 Finally, to reduce false positives, we implement a bug triage agent that identifies false alarms, enabling iterative refinement of the checkers.
-Figure 3. Overview of KNighter.
 3. Design
@@
 In the subsequent checker refinement phase, these valid checkers are deployed to scan the entire codebase for potential bugs.
-When bug reports are generated, a bug-triage agent evaluates them for false positives, and KNighter refines the checker accordingly.
-If the scan produces a manageable number of reports with a low false positive rate, KNighter presents the plausible checkers and the filtered reports as potential bugs for review.
+When bug reports are generated, a triage agent evaluates them for false positives, and KNighter refines the checker accordingly.
+If the scan produces a manageable number of reports with a low false positive rate, KNighter presents the plausible checkers and their filtered reports as potential bugs for review.
 Figure 4. Patch commit message.
@@
 Algorithm 1 Synthesize checkers with input patch.
+(a) Prompt template for bug pattern analysis
+(b) Prompt template for plan synthesis.
+(c) Prompt template for report triage.
+Figure 5. Simplified prompt templates used by KNighter.
 3.1.1. Bug Pattern Analysis
@@
 By including the signatures and brief descriptions of these utility functions in the prompt, we enable LLMs to leverage them effectively during the planning process, simplifying the overall task.
-(a) Prompt template for bug pattern analysis
-(b) Prompt template for plan synthesis.
-Figure 5. Prompt templates used by KNighter.
-3.1.3. Implementation and Syntax Repair
+3.1.3. Analyzer Implementation and Syntax Repair
+Figure 6. Pre-defined checker template for CSA.
 After identifying the bug pattern and making the plan, we leverage an LLM-based agent to implement the corresponding checker.
@@
 This automated debugging pipeline ensures that the final checkers are both syntactically correct and compilable.
-Figure 6. Pre-defined checker template for CSA.
 3.1.4. Validation
@@
 More details are in § 4.
-Figure 7. A report labeled as FP by our triage agent.
 3.2. Checker Refinement
@@
 Second, to navigate the complexity of analysis and modification, we employ specialized LLM-based agents.
-A triage agent classifies each distilled report, focusing strictly on pattern alignment (i.e., does the report match the original input bug’s pattern?) rather than general code correctness.
+A triage agent classifies each distilled report, focusing strictly on alignment with the target bug pattern (rather than general code correctness); the prompt template is shown in Figure 5(c).
 If the triage agent identifies a report as a false positive, as exemplified in Figure 7, a dedicated refinement agent takes over.
@@
 The refinement agent then uses this information to adjust the checker’s logic, specifically enhancing its ability to handle constructs like unlikely(), thereby preventing this type of false positive in subsequent scans while ensuring it can still detect the original vulnerability.
+Figure 7. A report labeled as FP by our triage agent.
 A refined checker is accepted only if it satisfies two criteria: (1) it no longer generates warnings for the previously identified false positive cases, and (2) it maintains its validity by correctly differentiating between the original buggy and patched code versions.
@@
 Our synthesized checker did not incorporate these nuanced constraints, and the triage agent likewise failed to recognize the critical differences, ultimately leading to a false positive.
+Table 2. Newly detected bugs by KNighter.
 Total Confirmed Fixed Pending CVE
-KNighter 92 77 57 15 16
-(a) Bug summary statistics detected by KNighter.
-(b) Number of bugs in each type.
-(c) Number of bugs in each subsystem.
-(d) Number of bugs with different lifetimes.
-Figure 9. The details of newly detected bugs by KNighter. Subfigure (9(a)) summarizes overall status, while (9(b)), (9(c)), and (9(d)) show breakdowns by type, subsystem, and lifetime.
+KNighter 92 77 57 15 30
+(a) Number of bugs in each type.
+(b) Number of bugs in each subsystem.
+(c) Number of bugs with different lifetimes.
+(d) Number of bugs detected by each commit.
+Figure 9. Details of new bugs. Subfigures (9(a)), (9(b)), (9(c)), and (9(d)) show breakdowns by type, subsystem, lifetime, and commit.
 5.2. RQ2: Detected Bugs
@@
 To date, static analyzers synthesized by KNighter have identified 92 new bugs in the Linux kernel.
-As summarized in Figure 9(a), developer confirmation has been received for 77 of these bugs.
+As summarized in Table 2, developer confirmation has been received for 77 of these bugs.
 Among those confirmed, 57 have already been fixed.
 The remaining 15 bugs are currently awaiting developer review (calculated as Total - Confirmed).
-Notably, 16 of the discovered bugs have been assigned CVE numbers, showing their practical security impact.
+Notably, 30 of the discovered bugs have been assigned CVE numbers, showing their practical security impact.
+The checkers for bug detection originate from two sources: (i) the initial 61 manually collected commits used for evaluation in § 3.1 across diverse bug types (shown in Table 1), and (ii) an additional set of 100 commits automatically collected using keywords related to Null-Pointer-Dereference to further explore that specific bug class.
+Figure 9(a) and Figure 9(b) show the bug distribution from these sources—with the light blue for bugs from the initial evaluation set and the light purple for those from the auto-collected set.
+Bug types.
+Analysis of checkers from the manually collected commits reveals that KNighter can detect a diverse range of bug types (see Figure 9(a)).
+Null-Pointer-Dereference bugs are the most prevalent, highlighting KNighter ’s strength in this area.
+In response, we expanded our effort by automatically collecting commits related to Null-Pointer-Dereference, which yielded an additional 30 bugs.
+Bug location.
+As shown in Figure 9(b), the detected bugs span various Linux kernel subsystems.
+The majority appear in the drivers subsystem (67 out of 92), reflecting its large footprint in the kernel (bursey2024syzretrospector, ).
+Additionally, 10 and 7 bugs were identified in the sound and net subsystems, respectively.
+Notably, 2 bugs were found in the samples directory—an area that provides example usage for kernel developers and where correctness is especially critical.
+Bug lifetime.
+Figure 9(c) illustrates the distribution of bug lifetimes.
+Notably, the average bug lifetime is 4.3 years, and 26 bugs had existed for over five years before detection.
+This indicates that the vulnerabilities uncovered by KNighter are difficult to detect and remain latent for extended periods, underscoring the effectiveness of our approach.
+Bug distribution over commits.
+Synthesized checkers from 39 bug-fix commits uncovered new bugs (2.4 each on average), with a long-tail skew (as shown in Figure 9(d)): five checkers each found five or more bugs.
+In general, checkers derived from recurring error patterns yield higher counts, while those from specialized fixes surface fewer, yet still valuable, findings.
+This suggests that KNighter can learn and propagate impactful patterns beyond their original contexts, producing a mix of broad-coverage and high-yield checkers.
+5.2.2. Case Study.
+Here are examples of vulnerabilities detected by KNighter.
 (a) Input Use-After-Free patch.
@@
 Figure 10. Example vulnerabilities detected by KNighter.
-The checkers for bug detection originate from two sources: (i) the initial 61 manually collected commits used for evaluation in § 3.1 across diverse bug types (shown in Table 1), and (ii) an additional set of 100 commits automatically collected using keywords related to Null-Pointer-Dereference to further explore that specific bug class.
-Figure 9(b) and Figure 9(c) show the bug distribution from these sources—with the light blue for bugs from the initial evaluation set and the light purple for those from the auto-collected set.
-Bug types.
-Analysis of checkers from the manually collected commits reveals that KNighter can find a wide range of bug types (in Figure 9(b)).
-Among these, Null-Pointer-Dereference bugs are the most prevalent, showing the KNighter’s strength in this area.
-In response, we expanded our effort by automatically collecting commits related to Null-Pointer-Dereference, which yielded an additional 30 bugs.
-Bug location.
-As shown in Figure 9(c), the detected bugs span various Linux kernel subsystems.
-The majority appear in the drivers subsystem (62 out of 92), reflecting its large footprint in the kernel (bursey2024syzretrospector, ).
-Additionally, 14 and 7 bugs were identified in the sound and net subsystems, respectively.
-Notably, 2 bugs were found in the samples directory—an area that provides example usage for kernel developers and where correctness is especially critical.
-Bug lifetime.
-Figure 9(d) illustrates the distribution of bug lifetimes.
-Notably, the average bug lifetime is 4.3 years, and 26 bugs had existed for over five years before detection.
-This indicates that the vulnerabilities uncovered by KNighter are difficult to detect and remain latent for extended periods, underscoring the effectiveness of our approach.
-5.2.2. Case Study.
-Here are examples of vulnerabilities detected by KNighter.
 CVE-2025-21715.
@@
 The absence of false negatives is particularly important, as it indicates the agent effectively prioritized all potentially true bugs in this set, minimizing the risk of overlooking genuine issues, even though the 22 false positives require further filtering.
+We also evaluated 5-way self-consistency via majority voting (wang2022self, ).
+For each report, we ran our triage agent five independent times and labeled the report as a “bug” only if the agent made that prediction in at least $t$ of the runs.
+Compared to our single-sample baseline, which identified 7 TPs and 22 FPs, majority voting did not offer a significant improvement.
+Using a threshold of $t=3$ kept the TP count at 7 but slightly increased FPs to 24.
+A stricter threshold of $t=4$ also resulted in 7 TPs while reducing FPs to 20.
+Ultimately, majority voting only slightly shifted the false positive count in this setting.
+There could be two potential reasons.
+For TPs, the agent likely identifies these bugs with high confidence, meaning a single run is sufficient to detect all of them.
+For FPs, the agent’s predictions are likely less confident and more inconsistent across different runs.
 5.4.2. Ablation Study
@@
 This highlights a crucial insight: successful checker synthesis demands more than general coding proficiency; accurate knowledge or inference of the target framework’s specific APIs and conventions (CSA in this case) is essential.
+6. Limitations and Discussion
+Limitations.
+KNighter faces challenges with highly complex bug patterns, particularly those involving state-machine reasoning, such as use-after-free, and concurrency issues that require analyzing multi-threaded code and locking schemes.
+Synthesizing precise checkers for these issues is difficult for today’s LLMs, as it requires a sophisticated understanding of a program’s temporal and interprocedural behavior.
+To address this, a promising future direction is to improve the analysis agent to automatically abstract complex bug reports into more formal state-machine representations.
+This would enable the synthesis of precise, state-aware checkers that can trace the specific sequences of operations leading to a bug.
+A second, complementary strategy is to focus on selectively identifying high-quality, canonical bug fixes.
+This involves developing methods to filter out patches that are overly complex or specific, allowing KNighter to learn from reusable, idiomatic repair patterns and improve the overall performance of checker synthesis.
+Generability.
+KNighter employs a general three-stage workflow: LLM-driven synthesis of checkers from bug-fix patches, patch-grounded validation, and triage/refinement to reduce false positives.
+While the Linux kernel served as an ideal and challenging initial target due to its scale, complexity, and importance, our approach is not fundamentally tied to it.
+The workflow’s applicability extends to any project that meets two criteria: a version history with bug-fix commits for learning, and a static analysis framework to serve as a synthesis target (e.g., Chromium (chromium, )).
+Furthermore, the implementation is not limited to C/C++.
+Although we demonstrated its use with the Clang Static Analyzer (clang-static-analysis, ), the pipeline can be readily adapted—via a small set of few-shot examples—to generate checkers or rules for other ecosystems (e.g., CodeQL (codeql, ) or Semgrep (semgrep, )).
 7. Related Work
+This section mainly discusses related work on both traditional and LLM-based static analysis, contextualizing the contributions of KNighter.
+While our work focuses on synthesizing static analyzers that inspect source code without execution, it is worth noting that checker synthesis has also been explored for generating dynamic runtime checks (t2c, ; lou2022demystifying, ; TrainCheckOSDI2025, ).
+These approaches are fundamentally different from KNighter.
+Dynamic checkers typically learn rules from test executions (t2c, ) to detect bugs based on concrete runtime states, such as capturing semantic “grey-failures”.
+In contrast, KNighter learns from source code patches to create static checkers that can cover all potential execution paths, including corner cases rarely triggered during program execution.
 7.1. Traditional Static Analysis
@@
 While often effective for their targeted bug classes, these approaches typically require extensive human expertise for analysis design and implementation, limiting their scalability and adaptability to newly emerging bug patterns in the rapidly evolving kernel.
+Deviation-based specification inference.
+Methods such as (engler2001bugs, ; min2015cross, ; ahmadi2021finding, ) infer specifications by assuming the majority of uses are correct and flagging deviations.
+While effective in some settings, classic systems rely on a small, fixed set of rule templates plus probabilistic clustering/ranking, which constrains bug-pattern coverage and—without strong semantic post-checks or refinement—often yields higher FPR (engler2001bugs, ; min2015cross, ).
+FICS (ahmadi2021finding, ) reduces template dependence via ML grouping of functionally similar code, but still rests on the majority-correct assumption and similarity heuristics, which can miscluster code and introduce noise.
 Patch-based specification inference.
@@
 Our synthesis approach.
-In contrast to the above, our work, KNighter, uses LLMs to synthesize the static analyzer itself directly from historical patch information.
+In contrast to the above, KNighter uses LLMs to synthesize the static analyzer itself directly from historical patch information.
 Instead of relying on predefined rules or inferring specifications for existing checkers, KNighter learns both the bug pattern and the corresponding detection logic, generating a new checker automatically.
@@
 Some techniques use LLMs to assist existing tools by automating tasks previously requiring manual effort.
-For example, LLMs can infer taint specifications for external APIs (IRIS (li2024llm, )), generate post-constraints to prune analysis paths (LLift (li2024enhancing, )), or infer resource-handling intentions (InferROI (wang2024resourceleak, )).
-While helpful, these methods still fundamentally depend on a substantial, human-developed analyzer core.
-This can limit their ability to easily adapt or generalize to detect novel bug types beyond those originally targeted by the core analyzer’s design.
+For example, LLMs can infer taint specifications for external APIs (IRIS (li2024llm, ), Artemis (ji2025artemis, )), generate post-constraints to prune analysis paths (LLift (li2024enhancing, )), or infer resource-handling intentions (InferROI (wang2024resourceleak, )).
+While helpful, these methods still fundamentally depend on a substantial, human-developed analyzer core, e.g., the manually defined rules for taint propagation (ji2025artemis, ).
+This can limit their ability to easily generalize to detect novel bug types beyond those originally targeted by the core analyzer’s design.
 Direct code analysis with LLMs.
@@
 Unlike LLM-augmented tools, this minimizes reliance on pre-existing manual analyzer development, enhancing adaptability for diverse bug types.
-Unlike direct LLM analysis, KNighter generates efficient, reusable static checkers, avoiding the high costs and scalability issues of scanning massive systems directly with LLMs.
-This makes KNighter a practical approach for applying LLM intelligence to large-scale static analysis, complementing other methods by automatically creating analyzers for new bug patterns.
-Checkers generated by KNighter could also potentially benefit from integration with LLM-augmented techniques.
+Unlike direct LLM analysis, KNighter generates efficient, reusable static checkers, avoiding the high costs and scalability issues of scanning massive codebases directly with LLMs.
+A very recent concurrent work, MoCQ (mocq, ), also explores checker/query synthesis, but it focuses on general bug patterns and relies on manual examples for validation.
+KNighter distinguishes itself not only by automatically inferring specific, nuanced patterns from patches but also by incorporating a closed-loop triage-refinement pipeline to iteratively improve checker precision.
+This fully automated refinement process makes our approach more scalable for detecting complex defects in system software, establishing it as a practical paradigm for applying LLM intelligence to large-scale static analysis.
 8. Conclusion
@@
 By synthesizing static analyzers rather than directly analyzing code, KNighter bridges the gap between LLMs’ reasoning capabilities and the practical constraints of analyzing massive systems.
-KNighter’s practical impact is shown by the discovery of 92 new, long-latent bugs in the Linux kernel, with 77 confirmed, 57 fixed, and 16 CVE assigned.
+KNighter’s practical impact is shown by the discovery of 92 new, long-latent bugs in the Linux kernel, with 77 confirmed, 57 fixed, and 30 CVE assigned.
 Looking forward, KNighter opens new possibilities for scalable LLM-based static analysis.
@@
 By leveraging LLMs to synthesize tools rather than perform analysis directly, we establish a scalable, reliable, and traceable paradigm for utilizing AI in critical software security applications.
+Acknowledgments
+We are grateful to the anonymous reviewers for their valuable feedback that helped to improve this paper. This work was partially supported by NSF grant CCF-2131943.